### v1

# Differentially private and decentralized randomized power method

Abstract
The randomized power method has gained significant interest due to its simplicity and efficient handling of large-scale spectral analysis and recommendation tasks. As modern datasets contain sensitive private information, we need to give formal guarantees on the possible privacy leaks caused by this method. This paper focuses on enhancing privacy-preserving variants of the method. We propose a strategy to reduce the variance of the noise introduced to achieve Differential Privacy (DP). We also adapt the method to a decentralized framework with a low computational and communication overhead, while preserving the accuracy. We leverage Secure Aggregation (a form of Multi-Party Computation) to allow the algorithm to perform computations using data distributed among multiple users or devices, without revealing individual data. We show that it is possible to use a noise scale in the decentralized setting that is similar to the one in the centralized setting. We improve upon existing convergence bounds for both the centralized and decentralized versions. The proposed method is especially relevant for decentralized applications such as distributed recommender systems, where privacy concerns are paramount.
Machine Learning, ICML
1 Introduction
The randomized power method has attracted significant attention due to its simplicity and its efficiency in solving large-scale linear algebra problems in spectral analysis or recommendation. It is in particular useful for computing the principal singular vectors of a matrix. It has a straightforward implementation which runs with low computational overhead (nearly linear complexity), making it particularly suitable for analyzing large modern datasets.
In large-scale machine learning systems, protecting user privacy is paramount. Datasets often contain sensitive personal or organizational information, so privacy has become a critical requirement. The standard randomized power method does not inherently provide privacy guarantees. While its output might seem less sensitive than the input data, there is no formal guarantee that private information embedded in the data cannot be inferred.
To address and quantify privacy leakages, Differential Privacy (DP) has emerged as a powerful mathematical framework that provides formal guarantees and helps to mitigate potential privacy leaks of an algorithm. DP ensures that the output of an algorithm reveals little about any individual record in the input, even in the worst case. Several works have attempted to apply DP to the randomized power method. For instance, (Hardt & Price, 2014) and (Balcan et al., 2016) introduced differentially private versions of the method, and (Guo et al., 2021; Wang & Xu, 2020) explores federated approaches to private power iteration. Other close works explore centralized differentially private versions of PCA (Liu et al., 2022) and federated (Wang & Xu, 2020; Briguglio et al., 2023) and give optimal convergence bounds given certain assumptions.
Despite these advancements, existing approaches suffer from several limitations. First, their performance heavily depends on the number of singular vectors being computed (Hardt & Price, 2014; Balcan et al., 2016; Guo et al., 2021; Liu et al., 2022; Wang & Xu, 2020), which impacts both utility and privacy guarantees. Second, they are primarily designed for centralized settings (Hardt & Price, 2014; Liu et al., 2022), where a trusted curator is assumed to hold the data. Moreover, some methods (Liu et al., 2022) make strong assumptions about the underlying data distribution (i.e. that the data is sub-Gaussian) which makes it harder to use these methods in practice. Some federated versions (Briguglio et al., 2023) claim to guarantee privacy by use of the federated setting, although it has been shown that decentralization does not offer privacy by design (Geiping et al., 2020). Finally, no fully decentralized versions exist to our knowledge. The previously proposed federated versions (Balcan et al., 2016; Guo et al., 2021; Wang & Xu, 2020) use public channel communication and ignore advances in Multi-Party-Computation (MPC). This makes them unsuitable for decentralized environments, such as distributed recommender systems and social networks, where data is partitioned across users and/or devices and communications are restricted to a predefined communication graph.
1.1 Contributions
The core focus of this article is to introduce improved privacy-preserving randomized power methods and to extend them to a decentralized setting, while maintaining computational efficiency and performance. We propose new convergence bounds for the (decentralized) privacy-preserving power method; these show that our proposed algorithm has a weaker dependence on the number of singular vectors that are computed.
We propose an improved differentially private version of the randomized power method, with tighter convergence bounds, reducing the noise required to achieve privacy and therefore improving the accuracy of the method. In the process of deriving these bounds, we present a new, more straightforward proof of a privacy result for the differentially private randomized power method. [^1] This proof addresses some minor mistakes in a privacy proof from (Hardt & Roth, 2012, 2013) that have been reproduced in several follow-up works (Hardt & Price, 2014; Balcan et al., 2016). The modified result allows us to use a wider range of privacy parameters.
Finally, we propose decentralized versions of the algorithms. These are particularly relevant in contexts where data is distributed across multiple users or devices, as is often the case in recommender systems. To address the privacy concerns of the users, without forcing them to rely on a trusted curator, we propose a federated privacy-preserving version of the method using Secure Aggregation (Bell et al., 2020), a form of Multi-Party Computation (MPC) and a fully decentralized version using GOPA (GOssip noise for Private Averaging) (Sabater et al., 2022). This allows users to keep their data local while contributing to a privacy-preserving global computation. Our decentralized approach preserves the effectiveness of the centralized version of the randomized power method while incorporating the privacy advantages of local differential privacy (DP). Notably, by employing Secure Aggregation, we are able to set the noise scale comparable to that of central DP, enhancing accuracy. This makes the proposed method a promising candidate for decentralized and privacy-conscious applications.
Our contributions can be summarized as follows:
  1. 1.
    We introduce a novel privacy-preserving randomized power method that improves upon existing convergence bounds, with a reduced sensitivity to the number of singular vectors.
  2. 2.
    We extend the method to a decentralized setting, utilizing Secure Aggregation to ensure privacy in a distributed environment without the need for a trusted curator.
  3. 3.
    We provide new privacy proofs that correct errors in previous works, ensuring rigorous privacy guarantees.
  4. 4.
    We propose a new adjacency model to guarantee the differential privacy of our method, to open the door to new and more realistic use cases.
  5. 5.
    We present both runtime-dependent and runtime-independent bounds for the proposed methods, demonstrating their utility through theoretical analysis and empirical validation.
2 Background Material
2.1 Notation:
Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a real-valued positive semi-definite matrix, and let $k$ denote the target rank, where typically $k\ll n$. We also define $\eta$ as a small positive scalar that represents the precision of the approximation.
The eigenvalue decomposition of $\bm{A}$ is given by:
$\bm{A}=\bm{U}\bm{\Lambda}\bm{U}^{\top},$ (1)
where $\bm{U}$ is the matrix of eigenvectors of $\bm{A}$, and $\bm{\Lambda}$ is a diagonal matrix containing the corresponding eigenvalues on its diagonal.
Let $\mathcal{N}(\mu,\sigma^{2})^{n\times p}$ be a ($n\times p$) random matrix of which each element is a random variable independent and identically distributed (i.i.d.) according to a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$.
For any matrix $\bm{X}$, let its element-wise maximum norm be $\|\bm{X}\|_{\max}=\max_{i,j}\|\bm{X}_{ij}\|$, where $\bm{X}_{ij}$ is the $(i,j)$-th entry of matrix $\bm{X}$. We denote by $\bm{X}_{j:}$ the $j$-th row of $\bm{X}$.
We also define two coherence measures for the matrix $\bm{A}$:
  • The $\mu_{0}$-coherence of $\bm{A}$ is the maximum absolute value of its eigenvector entries, defined as $\mu_{0}(\bm{A})=\|\bm{U}\|_{\max}=\max_{i,j}|\bm{U}_{ij}|$.
  • The $\mu_{1}$-coherence of $\bm{A}$ is the maximum row sum of the absolute values of its eigenvector entries, defined as $\mu_{1}(\bm{A})=\|\bm{U}\|_{\infty}=\max_{i}(\sum_{j}|\bm{U}_{ij}|)$.
Finally, we will also use the QR decomposition of matrix obtained using the Gram-Schmidt procedure. Given a matrix $\bm{X}\in\mathbb{R}^{n\times p}$, the QR decomposition factorizes it as $\bm{X}=\bm{Q}\bm{R}$, where $\bm{Q}$ is an orthonormal matrix (i.e., $\bm{Q}^{\top}\bm{Q}=\bm{I}$) and $\bm{R}$ is an upper triangular matrix.
2.2 Differential privacy:
2.2.1 Approximate differential privacy:
A randomized algorithm $\mathcal{M}$ provides ($\epsilon,\delta$)-differential privacy if for all datasets $D_{1}$ and $D_{2}$ differing on at most one element (neighboring datasets), and for all measurable subsets $S\subseteq\text{Range}(\mathcal{M})$, the following holds:
$\Pr[\mathcal{M}(D_{1})\in S]\leq e^{\epsilon}\Pr[\mathcal{M}(D_{2})\in S]+\delta,$ (2)
where $\epsilon$ is a small positive parameter representing the privacy loss (smaller values indicate stronger privacy guarantees), and $\delta$ is a small probability representing the chance that the privacy guarantee may fail, allowing for a small probability of privacy breach.
2.2.2 $\ell_{2}$-sensitivity:
The $\ell_{2}$-sensitivity of a function $f$, denoted as $\Delta_{2}(f)$ or $\Delta_{2}$ for brevity, measures the maximum change in the output of $f$ when its input changes by at most one element (for neighboring datasets $D_{1}\sim D_{2}$). It is defined as:
$\Delta_{2}=\max_{D_{1}\sim D_{2}}\|f(D_{1})-f(D_{2})\|_{2},$ (3)
where $\|\cdot\|_{2}$ represents the $\ell_{2}$-norm. This sensitivity quantifies the stability of the function under small changes in the input data and allows us to calibrate the noise needed to ensure differential privacy. The notion of adjacency of the datasets is in general application-dependent. Its granularity reflects what one wants to protect by using Differential Privacy. In (Hardt & Price, 2014; Balcan et al., 2016), it is defined as a magnitude 1 one-element change in the matrix of interest. In the context of recommender system, this would mean that we want to protect individual user-item interactions.
2.2.3 Zero-Concentrated Differential Privacy (zCDP):
A randomized algorithm $\mathcal{M}$ is said to satisfy $\rho$-zero-Concentrated Differential Privacy (zCDP) if for all neighboring datasets $D_{1}$ and $D_{2}$, and for all $\alpha\in(1,\infty)$, the following holds:
$D_{\alpha}(\mathcal{M}(D_{1})\|\mathcal{M}(D_{2}))\leq\rho\alpha\,,$ (4)
where $D_{\alpha}$ is the Rényi divergence of order $\alpha$ and $\rho$ is a positive parameter controlling the trade-off between privacy and accuracy (smaller values of $\rho$ imply stronger privacy guarantees). zCDP ensures that the privacy loss is tightly bounded in terms of the Rényi divergence, providing a more refined privacy analysis compared to pure or approximate differential privacy. Additionally, zCDP composes more easily than other differential privacy definitions and will be useful for our analysis.
The following lemma, introduced by Bun & Steinke (2016), specifies how the addition of Gaussian noise can be used to construct a randomized algorithm that satisfies zCDP.
Lemma 2.1 (Gaussian Mechanism (Proposition 1.6 (Bun & Steinke, 2016))).
Let $f:X^{n}\to\mathbb{R}$ be a sensitivity-$\Delta$ function. Consider the mechanism (randomized algorithm) $M:X^{n}\to\mathbb{R}$, defined as $M(x)=f(x)+Z_{x}$ where for each input $x$, $Z_{x}$ is independently drawn from $\mathcal{N}(0,\sigma^{2})$. Then $M$ satisfies $\left(\frac{\Delta^{2}}{2\sigma^{2}}\right)$-zCDP.
The next lemma, which is a generalized version of a result presented by Bun & Steinke (2016), explains how a randomized algorithm, constructed by recursively composing a sequence of zCDP-satisying randomized algorithms, also satisfies zCDP.
Lemma 2.2 (Adaptive composition (Generalization from Lemma 2.3 (Bun & Steinke, 2016))).
Let $M_{1}:X^{n}\to Y_{1}$, $M_{2}:X^{n}\times Y_{1}\to Y_{2}$, …, $M_{L}:X^{n}\times Y_{1}\times\dots\times Y_{L-1}\to Y_{L}$ be randomized algorithms. Suppose $M_{i}$ satisfies $\rho_{i}$-zCDP as a function of its first argument for each $i=1,2,\dots,L$. Let $M^{\prime\prime}:X^{n}\to Y_{L}$, constructed recursively by:
$M^{\prime\prime}(x)=M_{L}(x,M_{L-1}(x,\dots,M_{2}(x,M_{1}(x))\dots)).$ (5)
Then $M^{\prime\prime}$ satisfies $(\sum_{i=1}^{L}\rho_{i})$-zCDP.
2.3 Privacy preserving power iteration:
Let $k$ be a target rank and $\eta$ a small real controlling the approximation precision. The aim of the randomized power method is to find a matrix $\bm{Q}\in\mathbb{R}^{n\times k}$ such that $\|\bm{U}-\bm{Q}\bm{Q}^{\top}\bm{U}\|\leq\eta$. In other words, we want $\bm{Q}$ to approximate the $k$ principal eigenvectors of $\bm{A}$. The aim of the privacy-preserving randomized power method is to compute $\bm{Q}$ without leaking sensitive information embedded in $\bm{A}$. In (Hardt & Price, 2014; Balcan et al., 2016), a sensitive change is modeled as a single-element change of maximum Frobenius norm 1 in $\bm{A}$. According to (Hardt & Price, 2014; Balcan et al., 2016), given approximate-DP parameters ($\epsilon,\delta$) we can compute $\bm{Q}$ using Algorithm 1 with $\Delta_{l}=\Delta_{l}^{a}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$.
1:  Input: Matrix $\bm{A}\in\mathbb{R}^{n\times n}$, number of iterations $L$, target rank $k$, iteration rank $p\geq k$, privacy parameters $\epsilon,\delta$.
2:  Output: approximated eigen-space $\bm{X}^{L}\in\mathbb{R}^{n\times p}$, with orthonormal columns.
3:  Initialization: orthonormal $\bm{X}^{0}\in\mathbb{R}^{n\times p}$ by QR decomposition on a random Gaussian matrix $\bm{G}_{0}$; noise variance parameter $\sigma=\epsilon^{-1}\sqrt{{4L\log(1/\delta)}}$;
4:  for $\ell=1$ to $L$ do
5:      1. Compute $\bm{Y}_{\ell}=\bm{A}\bm{X}^{\ell-1}+\bm{G}_{\ell}$ with $\bm{G}_{\ell}\sim\mathcal{N}(0,\sigma_{l}^{2}={\Delta}_{l}^{2}\cdot\sigma^{2})^{n\times p}$2. Compute QR factorization $\bm{Y}_{\ell}=\bm{X}^{\ell}\bm{R}_{\ell}$.
6:  end for
Algorithm 1 Privacy-preserving randomized power method
In recent works on the privacy-preserving power iteration (Hardt & Price, 2014; Balcan et al., 2016; Guo et al., 2021), the standard deviation of the noise related to Differential Privacy (DP) at iteration $l$ is:
$\sigma_{l}$ $=\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$ (6)
$=\Delta_{l}^{a}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)},$ (7)
as defined in Fig. 3 of (Hardt & Price, 2014), Alg. 2 of (Balcan et al., 2016) and Theorem 6 of (Guo et al., 2021). $\Delta_{l}^{a}$ is an $\ell_{2}$-sensitivity bound of the change in $\bm{A}\bm{X}^{\ell-1}$ given a change of magnitude at most 1 in $\bm{A}$ on only one element ($\Delta_{l}\leq\Delta_{l}^{a}$). Our first contribution comes from the fact that we can find a significantly tighter bound on $\Delta_{l}$.
2.4 Existing convergence bounds
To our knowledge, the strongest bound for the randomized power method as defined in Algorithm 1 with $\Delta_{l}=\Delta_{l}^{a}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$ is given in (Balcan et al., 2016). In Theorem 2.4, we present a reduction of their Theorem 3.1. The theorem, and several others we present relies on the verification of the conditions in Assumption 2.3.
Assumption 2.3.
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$, and iteration rank $p$, with $k\leq q\leq p$. Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues. Let us fix:
$\eta=O\left(\frac{\sigma_{q}}{\sigma_{k}}\cdot\min\left\{\frac{1}{\log\left(\frac{\sigma_{k}}{\sigma_{q}}\right)},\frac{1}{\log\left(\tau d\right)}\right\}\right).$ (8)
Assume that at every iteration $l$ of Algorithm 1, $\bm{G}_{\ell}$ satisfies, for some constant $\tau>0$:
$\|\bm{G}_{\ell}\|_{2}$ $=O\left(\eta(\sigma_{k}-\sigma_{q+1})\right)\,,\quad\text{and}$ (9)
$\quad\|\bm{U}_{q}^{\top}\bm{G}_{\ell}\|_{2}$ $=O\left(\eta\left(\sigma_{k}-\sigma_{q+1}\right)\frac{\sqrt{p}-\sqrt{q-1}}{\tau\sqrt{d}}\right)\,.$ (10)
Theorem 2.4 (Private Power Method (PPM) (Balcan et al., 2016)).
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$, and iteration rank $p$, with $2q\leq p\leq d$. Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$. Let $\epsilon,\delta\in(0,1)$ be the differential privacy parameters. Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues. Then if Assumption 2.3 holds, Algorithm 1 with $\Delta_{l}=\Delta_{l}^{a}$ gives us the following runtime-dependent convergence bound:
$\eta=O\left(\frac{\epsilon^{-1}\max_{l}{(\|\bm{X}_{l}\|_{\max})}\sqrt{pLd\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right)$
and also the following runtime-independent convergence bound
$\eta=O\left(\frac{\epsilon^{-1}\|\bm{U}\|_{\max}\sqrt{pLd\log(1/\delta)\log(d)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$
3 Analysis of Runtime-dependent Bounds
3.1 Proposed runtime-dependent bounds
3.1.1 Adjacency notion:
In (Hardt & Price, 2014; Balcan et al., 2016), a matrix $\bm{A}^{\prime}$ is denoted as adjacent to $\bm{A}$ ($\bm{A}\sim\bm{A}^{\prime}$) when $\bm{A}^{\prime}$ has a change of a single element of Frobenius norm at most $1$ compared to $\bm{A}$. In other words, we can write:
$\bm{A}^{\prime}=\bm{A}+c\cdot\bm{e}_{i}\cdot\bm{e}_{j}^{\top},$ (11)
with $\bm{e}_{i},\bm{e}_{j}\in\mathbb{R}^{d}$ being the canonical basis vector of dimension $d$. Without loss of generality, we have $0\leq i,j<n$ (the change of magnitude $c\leq 1$ can happen on any element $\bm{A}_{ij}$). In terms of differential privacy, this concept of adjacency allows us to protect element-wise updates to $\bm{A}$. We however note that as $\bm{A}$ and $\bm{A}^{\prime}$ are real and positive semi-definite, the change should also need to be on the diagonal of $\bm{A}$. We propose a new, more general, notion of adjacency to allow for other types of updates:
$\bm{A}^{\prime}=\bm{A}+\bm{C},$ (12)
with $\bm{C}$ a symmetric matrix representing the update, subject to $\sqrt{\sum_{i}\|C_{i:}\|_{1}^{2}}\leq 1$. The proposed adjacency notion is strictly more general than the previous adjacency notion (11). Indeed, $\bm{C}=c\cdot\bm{e}_{i}\cdot\bm{e}_{j}^{\top}$ is a special case. $\bm{C}$ can also for instance be non-zero on the diagonal or antidiagonal, or be another symmetric matrix.
3.1.2 Improved sensitivity bound:
The previous sensitivity bound $\Delta_{l}^{a}$ (used in (Balcan et al., 2016; Hardt & Price, 2014)) was found by using a straightforward extrapolation from the sensitivity of the change in $\bm{A}{\bm{X}^{\ell-1}}$ when ${\bm{X}^{\ell-1}}\in\mathbb{R}^{n\times 1}$ to when ${\bm{X}^{\ell-1}}\in\mathbb{R}^{n\times p}$. We can however find a tighter bound on $\Delta_{l}$ if we directly analyse the change in ${\bm{X}^{\ell-1}}\in\mathbb{R}^{n\times p}$. We give the result of our improved analysis using our proposed adjacency notion (12) in Theorem 3.1 and provide its proof in Appendix A.1.
Theorem 3.1 (Improved Sensitivity Bound).
Let $\Delta_{l}=\|\bm{A}^{\prime}\bm{X}^{l}-\bm{A}\bm{X}^{l}\|_{F}$ where $\bm{A}^{\prime}$ is defined using (12). Then $\Delta_{l}$ can be upped bounded as follows:
$\begin{split}\Delta_{l}\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{2}\triangleq\Delta_{l}^{b}\,.\end{split}$ (13)
3.1.3 Note on related privacy proofs:
Several Differential Privacy (DP) proofs for the private randomized power method (PPM) have been developed in prior works, for instance those by (Hardt & Price, 2014), (Balcan et al., 2016), (Hardt & Roth, 2013), (Hardt & Roth, 2012). These papers mostly rely on the same proofs to establish that the PPM satisfies ($\epsilon$, $\delta$)-DP. Specifically, (Balcan et al., 2016) references the privacy proof from (Hardt & Price, 2014), which builds upon the privacy proof from (Hardt & Roth, 2013). The proof therein is also closely related to the one of (Hardt & Roth, 2012).
However, both (Hardt & Roth, 2013) (Theorem 2.4) and (Hardt & Roth, 2012) (Theorem 2.4) contain errors in their proposed composition rules, where a comparison sign is mistakenly flipped. This error could potentially cause the privacy parameter $\epsilon$ in the proposed mechanism to be arbitrarily small, providing no privacy guarantee at all. The original composition rule is presented in Theorem III.3 of (Dwork et al., 2010). Moreover, Lemma 3.4 of (Hardt & Roth, 2013) misuses Theorem 2.4. Indeed, they claim that their algorithm satisfies $(\epsilon^{\prime},\delta)$-DP at each iteration. By their Theorem 2.4, then the algorithm overall should satisfy $(\epsilon^{\prime},k\delta+\delta^{\prime})$-DP where $\delta^{\prime}>0$, and not $(\epsilon^{\prime},\delta)$-DP as claimed.
3.1.4 Privacy proof:
To address these issues, we propose a new proof of the Differential Privacy guarantees for our algorithm in Theorem 3.2, whose proof is in Appendix A.2. This proof clarifies the ambiguities or errors present in the previous works, and ensures that our privacy guarantees are met.
Theorem 3.2 (Privacy proof for the PPM).
Assuming $\delta\in(0,1)$ and $\epsilon>0$ such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$, the Private Power Method with $\Delta_{l}=\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ satisfies ($\epsilon,\delta$)-Differential Privacy.
3.1.5 Improved convergence bound:
Theorem 3.3 (Bound for the noisy power method (NPM) (Balcan et al., 2016)).
Let reals $k\leq q\leq p$. Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of a positive semi-definite matrix $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues and fix $\eta=O\left(\frac{\sigma_{q}}{\sigma_{k}}\cdot\min\left\{\frac{1}{\log\left(\frac{\sigma_{k}}{\sigma_{q}}\right)},\frac{1}{\log\left(\tau d\right)}\right\}\right)$. If at every iteration $l$ of the NPM $\bm{G}_{\ell}$ satisfies Assumption 2.3, then after
$L=\Theta\left(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log\left(\frac{\tau d}{\eta}\right)\right).$
iterations, with probability at least $1-\tau^{-\Omega(p+1-q)}-e^{-\Omega(d)}$, we have
$\|(\bm{I}-\bm{X}^{L}{\bm{X}^{L}}^{\top})\bm{U}_{k}\|_{2}\leq\eta.$ (14)
Theorem 3.4 (Improved PPM with runtime dependent bound).
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$ and iteration rank $p$ with $2q\leq p\leq d$. Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$. Let $\delta\in(0,1)$ and $\epsilon>0$ be privacy parameters such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$
Then with approximate DP parameters ($\epsilon$,$\delta$) and assuming Assumption 2.3 is verified, Algorithm 1 with $\Delta_{l}=\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ gives us the following runtime-dependent convergence bound:
$\eta=O\left(\frac{\epsilon^{-1}\max_{i,l}\|\bm{X}_{i:}^{l}\|_{2}\sqrt{Ld\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$ (15)
Note: This bound is at least as tight as the one proposed in (Balcan et al., 2016). Indeed the following bound holds:
[…]
which is tight only if $\forall j,\|\bm{X}_{ij}^{l}\|=\max_{j}\|\bm{X}_{:j}^{l}\|$. and can be in practice much tighter depending on $\bm{A}$. Our privacy proof (given in Theorem 3.2) does not restrict us to have $\epsilon\leq 1$. We provide the proof in Appendix A.2.
4 Analysis of Runtime-independent bounds
The runtime-independent bound that we develop in this section and present in Theorem 4.1 is a combination of two bounds. The first applies for matrices with large coherences, and is independent of the matrix coherence parameter. The second applies to matrices with low $\mu_{1}$-coherence, but the bound is dependent on the $\mu_{1}$-coherence of the matrix. We also show that our method achieves a bound which is at least as tight as both proposed bounds.
4.1 Proposed runtime-independent convergence bound
The proposed runtime-dependent bound that we presented earlier (Theorem 3.4) involves $\max_{i}\|\bm{X}_{i:}\|_{2}$, which depends on values that change during the execution of the algorithm. Similarly to (Hardt & Price, 2014; Hardt & Roth, 2013), we now propose a runtime-independent convergence bound for our method. This bound, presented in Theorem 4.1, is a combination of two bounds, one that is suited for matrices with high $\mu_{1}(\bm{A})$, and another that applies for matrices with small $\mu_{1}(\bm{A})$. We provide its proof in Appendix A.3. We note that for a matrix $\bm{A}$, the proposed bound is tighter than the one proposed in (Balcan et al., 2016) when $\min(1,\mu_{1}(\bm{A}))\leq\sqrt{p\cdot\log({d})}\cdot\mu_{0}(\bm{A})$. This is likely to happen in practice. We highlight this fact in Section 5.1.
Theorem 4.1.
Improved PPM with runtime independent bound
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$ and iteration rank $p$ with $2q\leq p\leq d$. Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$. Let $\delta\in(0,1)$ and $\epsilon>0$ be privacy parameters such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$ Then with approximate DP parameters ($\epsilon$,$\delta$) and assuming Assumption 2.3 is verified, Algorithm 1 gives us the following runtime-independent convergence bound:
$\eta=O\left(\frac{\epsilon^{-1}\cdot\min(1,\mu_{1}(\bm{A}))\cdot\sqrt{Ld\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$ (17)
5 Empirical comparison of the proposed bounds
In previous sections, we proposed both runtime-dependent and runtime-independent bounds of convergence for our algorithm. The runtime-dependent bound is dependent on the matrix of interest $\bm{A}$ and on the matrices $\bm{X}_{l}$ that are computed iteratively. The runtime-independent bound (4.1) is dependent only on the matrix of interest $\bm{A}$. We now provide a few examples to justify the usefulness of the proposed bounds.
5.1 Empirical comparison of the runtime-independent bounds
As explained in Section 4.1, the proposed runtime-independent bound is tighter than those proposed in (Hardt & Price, 2014; Balcan et al., 2016) when $\min(1,\mu_{1}(\bm{A}))\leq\mu_{0}(\bm{A})\sqrt{p\cdot\log({d})}$, where $\bm{U}$ denotes the eigenvectors of $\bm{A}$. We now provide an illustration of how the bounds compare for a practically relevant example, namely recommender systems.
Table 1: Statistics of datasets
Dataset Users Items Interactions $\mu_{0}$ $\mu_{1}$ $\Delta_{l}^{a}$ (existing) ${\Delta_{l}^{b}}$ (ours)
Gowalla 29,858 40,981 1,027,370 0.36 82.48 1.08 ×$\times\sqrt{p}$ 1
Yelp2018 31,668 38,048 1,561,406 0.40 60.42 1.29 ×$\times\sqrt{p}$ 1
Amazon-book 52,643 91,599 2,984,108 0.33 52.95 1.10 ×$\times\sqrt{p}$ 1
5.1.1 Recommender systems:
To compute several parameters of SOTA recommender system models, we need to compute leading eigenvectors of a matrix denoted as “the normalized item-item matrix”. This is equivalent as computing the right singular vectors of “the normalized adjacency matrix” (Shen et al., 2021). The most widely used datasets in recommender systems research are Gowalla (Cho et al., 2011), Yelp2018 (He et al., 2020) and Amazon-Book (Wang et al., 2019). Table 1 presents key statistics of these datasets, and compares the sensitivity bounds for the task of computing the leading eigenvectors of interest. As we can deduce from Table 1, the proposed bound allows us to converge to solutions with much smaller $\eta$ values, especially when the desired number of factors $p$ is large.
5.2 Empirical comparison of the runtime-dependent bounds
[…]
$r(k,n)=\frac{\mathbb{E}[\Delta_{0}^{a}]}{\mathbb{E}[\Delta_{0}^{b}]}.$ (18)
5.2.1 Theoretical approximation of $r(k,n)$:
We can find an asymptotic approximation of $r(k,n)$ using Approximation 5.1. The derivation is provided in Appendix A.4.
Approximation 5.1.
[…]
$\mathbb{E}[\Delta_{l}^{a}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(kn)\frac{k^{2}}{n^{2}}}+\frac{k\mu}{n}\right)$ (19)
$\mathbb{E}[\Delta_{l}^{b}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(n)\frac{k}{n^{2}}}+\frac{k\mu}{n}\right).$ (20)
5.2.2 Empirical approximation of $r(k,n)$:
Figure 1: Comparison of empirically (in blue) and theoretically (in green) estimated $r(k,n)$ ratios, with $n=8000$ and $k$ ranging between 64 and 4000, with a step of 64. Blue dots represent empirical ratios estimated using $t=5$ runs of the first step of our algorithm. The theoretical ratios are estimated using Theorem 5.1.
Figure 2: Comparison of empirically (in blue) and theoretically (in green) estimated $r(k,n)$ ratios, with $n=12000$ and $k$ ranging between 64 and 4000, with a step of 64. Blue dots represent empirical ratios estimated using $t=5$ runs of the first step of our algorithm. The theoretical ratios are estimated using Theorem 5.1.
We now want to compare theoretically and empirically how tight our proposed bound ($\Delta_{0}^{b}$) is compared to the previous one ($\Delta_{0}^{a}$). The scaling of the noise at the first iteration of the algorithm depends on the random initialization of the eigenvector matrix $\bm{X}_{0}$, which does not depend on the matrix of interest $\bm{A}$. We can therefore empirically estimate $\mathbb{E}[\Delta_{0}^{a}]$ and $\mathbb{E}[\Delta_{0}^{b}]$ on randomly initialized matrices $\bm{X}_{0}$ to compare the tightness of both bounds.
Experiment: $r(k,n)$ depends on the number of factors $k$ and on $n$, where ($n\times n$) is the size of $\bm{A}$. We therefore seek to compute it for multiple values of $n$ ($n\in\{8000,12000\}$) and a range of values for $k$ (between 64 and 4000, with steps of 64), for each value of $n$. We use Theorem 5.1 to approximate it theoretically. To measure it empirically, we compute $\Delta_{0}^{a}$ and $\Delta_{0}^{b}$ based on $\bm{X}_{0}$ in the first step of the algorithm (it is independent of the matrix of interest $\bm{A}$), for $t$ runs of the algorithm. We can then use the $t$ measures to estimate $\mathbb{E}[\Delta_{0}^{a}]$ and $\mathbb{E}[\Delta_{0}^{b}]$. We claim that $t=5$ is enough to see a general tendancy, as we have estimates for many ($k,n$) couples. We can also compare the empirical estimates to the proposed asymptotic approximations specified in Approximation 5.1.
We give our results in Figures 1 and 2. We can see empirically and theoretically that the proposed noise scaling $\Delta_{l}^{b}$ is much tighter than $\Delta_{l}^{a}$ at the first step of the algorithm. Our theoretical approximation is conservative and underestimates how much tighter the bound is initially, compared to what we observe in practice. The proposed bound is tighter by a multiplicative factor on the first step and can therefore drastically reduce the impact of the noise introduced by DP on the convergence of the Privacy Preserving power method, as the algorithm is usually run for very few steps ($L$ is usually in the range of $1$-$5$).
6 Decentralized version
Table 2: Overhead of using Secure Aggregation in our protocol
Client Server
Communication $O(log(s)+n^{2})$ $O(s\cdot log(s)+sn^{2})$
Computation $O(log^{2}(s)+n^{2}\cdot log(s))$ $O(s\cdot log^{2}(s)+sn^{2}\cdot log(s))$
In this section, we explore the scenario where the matrix $\bm{A}$ is distributed across multiple clients. Specifically, each client holds a matrix $\bm{A}^{(1)},\cdots,\bm{A}^{(s)}\in\mathbb{R}^{n\times n}$ such that $\sum_{i=1}^{s}\bm{A}^{(s)}=\bm{A}$. The goal is for the clients to collaboratively compute an orthonormal basis for the range of $\bm{A}$, similar to the centralized randomized power method, but without revealing their individual private matrices $\bm{A}^{(s)}$ to the server or other clients.
The randomized power method involves linear operations, which makes it well-suited for parallelization. An implementation of a private and distributed power method, using public channel communication between the clients and the server, was proposed in (Balcan et al., 2016). However, this approach relies on local Differential Privacy, since the data exchanged between the clients and the server can be observed by everyone.
To improve privacy while keeping the benefits of distributed computation, we propose using Secure Aggregation, a lightweight Multi-Party Computation (MPC) protocol. Secure Aggregation eliminates the need for local DP by allowing us to use distributed DP, which offers privacy guarantees similar to central DP. This approach has been extensively studied in the literature (Chen et al., 2021; Kairouz et al., 2021; Goryczka et al., 2013; Wei et al., 2024; Ghazi et al., 2019).
In distributed DP, each client adds noise to its own contributions before using the Secure Aggregation protocol. The noise is calibrated such that the sum of the noise across clients has a variance comparable to that used in Central DP, thus achieving similar privacy guarantees without requiring a trusted curator to aggregate the data.
We assume the “honest-but-curious” threat model, meaning that while the clients may attempt to learn additional information, they follow the protocol honestly. We also assume no corrupted or dropout users. For simplicity, this paper focuses on a basic version of distributed DP, neglecting effects of data quantization or errors introduced by modular arithmetic. We refer to (Kairouz et al., 2021) for an implementation of Distributed DP that takes these effects into account.
Finally, it is also possible to use a fully decentralized and differentially private protocol to perform averaging (Sabater et al., 2022). It is therefore straightforward to transform our protocol to avoid relying on a central server for Secure Aggregation, while keeping similar utility and communication costs.
6.1 Decentralized and private power method using distributed DP:
We give a federated version of the PPM in Algorithm 2. In this version, we reduce the variance of the noise by a multiplicative factor in O($s\cdot p\log({d})\cdot\|\bm{U}\|_{\infty}^{2}$) compared to (Balcan et al., 2016) due to the fact that we emulate Central DP with Secure Aggregation (thus using a secure channel) rather than using a public channel that only allows for local DP, and because we propose a generally tighter sensitivity bound. We show the equivalence between Algorithm 2 and Algorithm 1 in Theorem 6.2, and provide the proof in Appendix A.5. We first define the secure aggregation algorithm.
Definition 6.1.
We denote by $SecAgg({\bm{Y}_{\ell}^{(i)}},\{i|1{\leq}i{\leq}s\})$ the Secure Aggregation of locally held matrices $\bm{Y}_{\ell}^{(i)}$ over the set of users represented by indices $\{i|1{\leq}i{\leq}s\})$. It is equivalent to computing $\bm{Y}_{\ell}=\sum_{i=1}^{s}\bm{Y}_{\ell}^{(i)}$ over a secure channel.
  Input: distributed matrices $\bm{A}^{(1)},\cdots,\bm{A}^{(s)}\in\mathbb{R}^{n\times n}$, number of iterations $L$, target rank $k$, iteration rank $p\geq k$, private parameters $\epsilon,\delta$.
  Output: approximated eigen-space $\bm{X}^{L}\in\mathbb{R}^{n\times p}$, with orthonormal columns.
  Initialization: orthonormal $\bm{X}^{0}\in\mathbb{R}^{n\times p}$ by QR decomposition on a random Gaussian matrix $\bm{G}_{0}$; noise variance parameter $\nu=\epsilon^{-1}\sqrt{\frac{4L\log(1/\delta)}{s}}$;
  for $\ell=1$ to $L$ do
      1. The central node broadcasts $\bm{X}^{\ell-1}$ to all $s$ computing nodes;2. Computing node $i$ computes $\bm{Y}^{(i)}_{\ell}=\bm{A}^{(i)}\bm{X}^{\ell-1}+\bm{G}^{(i)}_{\ell}$ with $\bm{G}^{(i)}_{\ell}\sim\mathcal{N}(0,\Delta_{l}^{2}\nu^{2})^{n\times p}$;3. The central node computes with the clients $\bm{Y}_{\ell}=SecAgg({\bm{Y}_{\ell}^{(i)}},\{i|1\leq i\leq s\})$. 4. The central node computes QR factorization $\bm{Y}_{\ell}=\bm{X}^{\ell}\bm{R}_{\ell}$.
  end for
Algorithm 2 Federated private power method
Theorem 6.2 (Privacy and convergence of Algorithm 2).
Algorithm 2 is as privacy-preserving and has the same utility as its centralized version (Algorithm 1).
6.2 Communication and computation cost analysis:
We denote by $s$ the number of shards, such that $\sum_{i=1}^{s}\bm{A}^{(i)}=\bm{A}\in\mathbb{R}^{n\times n}$. Table 2 presents the communication and computations cost overhead introduced by the polylogarithmic version of Secure Aggregation that we employ (Bell et al., 2020).
7 Conclusion
We presented a differentially private and decentralized randomized power method that addresses privacy concerns in large-scale spectral analysis and recommendation systems. We introduced new convergence bounds, improving the accuracy of the method while ensuring privacy guarantees. By employing Secure Aggregation in a decentralized setting, we were able to reduce the noise introduced for privacy, leading to higher utility without compromising on privacy. The decentralized variant of the algorithm maintains the efficiency and privacy of the centralized version but adapts it for distributed environments. Our theoretical and empirical results demonstrate the utility and privacy advantages of our approach, suggesting that it offers a strong solution for real-world decentralized applications.
8 Limitations
The proposed method is effective in ensuring differential privacy and decentralization but still has limitations. The privacy guarantees rely on the honest-but-curious threat model and assume no collusion or malicious behavior from participants, which may not always be realistic in certain environments. In the future, it would also be interesting to analyze a private and decentralized version of the accelerated version of the randomized power method, which could potentially have a faster convergence.
Appendix A Proofs of Results
A.1 Proof of Theorem 3.1
Proof.
We note $\bm{A}^{\prime}$ a matrix adjacent to $\bm{A}$ using (12).
$\bm{A}^{\prime}=\bm{A}+\bm{C},$ (21)
with $\bm{C}$ a symmetric matrix representing the update, subject to $\sqrt{\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}}\leq 1$.
Then
$\begin{split}\Delta_{l}&=\|\bm{A}\bm{X}^{l}-\bm{A}^{\prime}\bm{X}^{l}\|_{F}\\ &=\|\bm{A}\bm{X}^{l}-(\bm{A}+\bm{C})\bm{X}^{l}\|_{F}\\ &=\|\bm{C}\bm{X}^{l}\|_{F}\\ &=\sqrt{\sum_{i}\|\bm{C}_{i:}\bm{X}^{l}\|_{F}^{2}}.\\ \end{split}$ (22)
and
$\begin{split}\|\bm{C}_{i:}\bm{X}^{l}\|_{F}&=\|\sum_{j}\bm{C}_{ij}\cdot\bm{e}_{j}^{\top}\cdot\bm{X}^{l}\|_{F}\\ &=\|\sum_{j}\bm{C}_{ij}\cdot\bm{X}^{l}_{j:}\|_{F}\\ &\leq\sum_{j}\|\bm{C}_{ij}\cdot\bm{X}^{l}_{j:}\|_{F}\\ &\leq\sum_{j}|\bm{C}_{ij}|\cdot\|\bm{X}^{l}_{j:}\|_{F}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\cdot\sum_{j}|\bm{C}_{ij}|.\end{split}$ (23)
By injecting (22) into (23), we have:
$\begin{split}\Delta_{l}&=\sqrt{\sum_{i}\|\bm{C}_{i:}\bm{X}^{l}\|_{F}^{2}}\\ &\leq\sqrt{\sum_{i}(\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\cdot\sum_{j}|\bm{C}_{ij}|)^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}(\cdot\sum_{j}|\bm{C}_{ij}|)^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\triangleq\Delta_{l}^{b}.\\ \end{split}$ (24)
Although the proposed update model (12) is more general than when $\bm{A}^{\prime}$ is defined using (11), the proposed bound $\Delta_{l}^{b}$ is also generally tighter than the bound proposed in (Hardt & Price, 2014; Balcan et al., 2016).
A.2 Proof of Theorem 3.2
Proof.
We can see lines 4-5 of Algorithm 1 as a sequential composition ($M$) of $L$ Gaussian Mechanisms. By Lemma 2.1, each mechanism $M_{i}$ satisfies ($\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$)-zCDP where $\Delta_{i}$ is the $\ell_{2}$-sensitivity of the function associated to mechanism $\Delta_{i}$ and $\sigma_{i}^{2}$ is the variance of the noise added with the Gaussian Mechanism. By Lemma 2.2, the composition of mechanisms $M=(M_{1},...,M_{i},...,M_{L})$ satisfies ($\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$)-zCDP. Let $\rho\triangleq\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$. By design of our algorithm, we have:
$\rho$ $=\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$ (25)
$\leq\sum_{i=1}^{L}\frac{(\Delta_{l}^{b})^{2}}{2\sigma_{i}^{2}}$ (26)
$=\sum_{i=1}^{L}\frac{1}{2\sigma^{2}}$ (27)
$=\sum_{i=1}^{L}\frac{\epsilon^{2}}{{8L\log(1/\delta)}}$ (28)
$=\frac{\epsilon^{2}}{{8\log(1/\delta)}}.$ (29)
By Proposition 1.3 of (Bun & Steinke, 2016), if M provides $\rho$-zCDP, then M is ($\rho+2\sqrt{\rho\log(1/\delta)},\delta)$-DP, $\forall\delta>0$. Let $\epsilon^{\prime}\triangleq\rho+2\sqrt{\rho\log(1/\delta)}$.
Then:
$\epsilon^{\prime}$ $=\rho+2\sqrt{\rho\log(1/\delta)}$ (31)
$\leq\frac{\epsilon^{2}}{{8\log(1/\delta)}}+2\sqrt{\frac{\epsilon^{2}}{{8\log(1/\delta)}}.\log(1/\delta)}$ (32)
$\leq\frac{\epsilon^{2}}{{8\log(1/\delta)}}+\frac{\epsilon}{2}.$ (33)
To satisfy ($\epsilon$, $\delta$)-DP, we need:
$\epsilon\geq\epsilon^{\prime}$ $\impliedby\epsilon\geq\frac{\epsilon^{2}}{{8\log(1/\delta)}}+\frac{\epsilon}{2}$ (35)
$\iff\frac{\epsilon}{2}\geq\frac{\epsilon^{2}}{{8\log(1/\delta)}}$ (36)
$\iff\epsilon\leq 4\log(1/\delta)$ (37)
$\iff\delta\leq\exp{(-\frac{\epsilon}{4})},$ (38)
which is a reasonable assumption, as in practice $\epsilon=O(1)$ and $\delta\ll\frac{1}{d}$ where $d$ is the number of records to protect. In our case, $d=n^{2}$ because we run the privacy-preserving power method on $\bm{A}\in\mathbb{R}^{n\times n}$. ∎
A.3 Proof of Theorem 3.4
Proof.
According to (Hardt & Price, 2014), if $\bm{G}_{l}\sim\mathcal{N}(0,\sigma_{l})^{d\times p}$, then with probability p=99/100 we have:
$\begin{split}&\max_{l=1}^{L}\|\bm{G}_{l}\|\leq\sigma_{l}\cdot\sqrt{d\cdot\log(L)}\,,\\ &\max_{l=1}^{L}\|\bm{U}^{\top}\bm{G}_{l}\|\leq\sigma_{l}\cdot\sqrt{p\cdot\log(L)}\,.\end{split}$ (40)
We can therefore satisfy the noise conditions of Theorem 3.3 with probability $p$ if we choose $\eta=\frac{\sigma_{l}\cdot\sqrt{d\cdot\log(L)}}{\sigma_{k}-\sigma_{q+1}}$, giving us:
$\|(\bm{I}-\bm{X}^{L}{\bm{X}^{L}}^{\top})\cdot\bm{U}_{k}\|\leq\frac{\sigma_{l}\cdot\sqrt{d\cdot\log(L)}}{\sigma_{k}-\sigma_{q+1}}\,.$ (41)
Which leads us to the following Theorem by injecting $\sigma_{l}=\Delta_{l}^{b}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$.
A.4 Derivation of Approximation 5.1
We provide here the derivation for Approximation 5.1:
Proof.
Let $\bm{M}\in\mathbb{R}^{n\times k}$ with i.i.d $\mathcal{N}(0,1)$-distributed entries. Let $\bm{M}=\bm{Q}\bm{R}$ be its QR factorization (by definition $\bm{i}sQ$ orthogonal and $\bm{R}$ upper triangular). By the Barlett decomposition theorem (Muirhead, 2009), we know that $\bm{Q}$ is a random matrix distributed uniformly in the Stiefel manifold $\mathbb{V}_{k,n}$. Then by Theorem 2.2.1 of (Chikuse, 2012), we know that a random matrix $\bm{Q}$ uniformly distributed on $\mathbb{V}_{k,n}$ can be expressed as $\bm{Q}=\bm{Z}(\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}}$ with $\bm{Z}$ another matrix with i.i.d $\mathcal{N}(0,1)$-distributed entries. We approximate $\bm{Z}^{\top}\bm{Z}$ as a diagonal matrix and remark that its diagonal elements are distributed as random chi-squared variables with $n$ degrees of liberty. For a matrix $\bm{A}$, we note as $\bm{A}^{\circ n}$ the elementwise Hadamard exponentiation. Then we have:
$\bm{Q}_{ij}^{2}$ $=(\bm{Z}(\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}})_{ij}^{2}$ (42)
$=(\bm{Z}^{\circ 2}((\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}})^{\circ 2})_{ij}$ (43)
$=(\bm{Z}^{\circ 2}(\bm{Z}^{\top}\bm{Z})^{-1})_{ij}.$ (44)
We remark that each element of $\bm{Z}^{\circ 2}$ is distributed as a chi-squared variable with one degree of freedom, therefore $n\cdot\bm{Q}_{ij}^{2}$ is distributed as an i.i.d $F(1,n)$ random variable by definition. For large $n$, we can approximate these $F$ variables as Gaussians with mean $\mu=\frac{n}{n-2}$ and variances $\sigma^{2}=\frac{2n^{2}(n-1)}{(n-2)^{2}(n-4)}$. Therefore $\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{\mu}{n},\frac{\sigma^{2}}{n^{2}})$ and:
$k\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{k\mu}{n},\frac{k^{2}\sigma^{2}}{n^{2}})$ $\sum_{j}\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{k\mu}{n},\frac{k\sigma^{2}}{n^{2}}).$ (45)
We can approximate the expectancy of the maximum (noted as m) of $d$ Gaussian variables distributed as $\mathcal{N}(\mu_{2},\sigma_{2}^{2})$ by $m=\sigma_{2}\sqrt{2\cdot\log(d)}+\mu_{2}$ by Lemma 2.3 of (Massart, 1896). We therefore get approximations of $\mathbb{E}[\max_{ij}k\bm{Q}_{ij}^{2}]$ and $\mathbb{E}[\max_{i}\|\bm{Q}_{i:}\|^{2}]$. We assume that the variances of $\Delta_{l}^{a}$ and $\Delta_{l}^{b}$ (as $n$ and $k$ are large) are small and therefore:
$\mathbb{E}[\Delta_{l}^{a}]^{2}\approx\mathbb{E}[(\Delta_{l}^{a})^{2}]=\sqrt{\frac{k^{2}\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(kn)}+\frac{k\mu}{n}$ (46)
$\mathbb{E}[\Delta_{l}^{b}]^{2}\approx\mathbb{E}[(\Delta_{l}^{b})^{2}]=\sqrt{\frac{k\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(n)}+\frac{k\mu}{n}.$ (47)
A.5 Proof of Theorem 4.1
We provide here the proof for Theorem 4.1: See 4.1
Proof.
First bound: We know by Section 2.4 of (Woodruff et al., 2014) that any row $\bm{v}$ of a matrix with orthonormal columns $\bm{Z}$ has squared norm at most 1. Therefore $\max_{i}\|\bm{X}_{i:}\|_{2}\leq 1$ because $\bm{X}$ has orthonormal columns by construction.
Second bound: Without loss of generality and to simplify notation, we note $\bm{X}$ to denote a matrix $\bm{X}_{l}$ computed during the PPM. As proposed in (Hardt & Price, 2014), we can write any column $x_{c}$ of $\bm{X}$ ($\bm{X}_{:c}$) as $\sum_{i=1}^{n}\alpha_{c}^{i}u_{i}$ where $u_{i}$ denotes the $i$-th eigenvector of $\bm{A}$, and $\alpha_{c}^{i}$ is a scalar. Therefore we have:
$<x_{c},x_{o}>$ $=<\sum_{i=1}^{n}\alpha_{c}^{i}u_{i},\sum_{j=1}^{n}\alpha_{o}^{i}u_{j}>$ (48)
$=\sum_{i=1}^{n}\sum_{j=1}^{n}\alpha_{c}^{i}\alpha_{o}^{i}<u_{i},u_{j}>$ (By bilinearity) (49)
$=\sum_{i=1}^{n}\alpha_{c}^{i}\alpha_{o}^{i}.$ (Orth. columns of $\bm{U}$) (50)
$\bm{X}$ has orthonormal columns, so if $c=o$, then $<x_{c},x_{o}>=\sum_{i=1}^{n}{(\alpha_{c}^{i})}^{2}=1$. Otherwise, we have $<x_{c},x_{o}>=\sum_{i=1}^{n}\alpha_{c}^{i}\alpha_{o}^{i}=0$. We can therefore define a matrix $\bm{B}$ with orthonormal columns such that $\bm{X}=\bm{U}\bm{B}$ and $\bm{B}_{jc}=\alpha_{c}^{j}$. We can therefore write any row $i$ of $\bm{X}$ as $\bm{X}_{i:}=\sum_{j=1}^{p}\bm{U}_{ij}\bm{B}_{j:}$. Therefore
$\|\bm{X}_{i:}\|_{2}$ $=\|\sum_{j=1}^{p}\bm{U}_{ij}\bm{B}_{j:}\|_{2}$ (51)
$\leq\sum_{j=1}^{p}\|\bm{U}_{ij}\bm{B}_{j:}\|_{2}$ (52)
$\leq\sum_{j=1}^{p}|\bm{U}_{ij}|\|\bm{B}_{j:}\|_{2}$ (53)
$\leq\sum_{j=1}^{p}|\bm{U}_{ij}|.$ (Orth. columns of $\bm{B}$) (54)
We then have $\max_{i}\|\bm{X}_{i:}\|_{2}\leq\max_{i}(\sum_{j=1}^{p}|\bm{U}_{ij}|)=\|\bm{U}\|_{\infty}=\mu_{1}(\bm{A})$.
We proposed two bounds for $\|\bm{X}_{i:}\|_{2}$, which are useful in two different cases. We can however note that:
$\|\bm{X}_{i:}\|_{2}\leq 1\text{ and }\|\bm{X}_{i:}\|_{2}\leq\mu_{1}(\bm{A})$ (55)
$\Rightarrow\|\bm{X}_{i:}\|_{2}\leq\min(1,\mu_{1}(\bm{A})),$ (56)
which leads us to Theorem 4.1 giving us a runtime-independent bound.
A.6 Proof of Theorem 6.2
Proof.
It is straightforward to see that if steps 2 and 3 of Algorithm 2 are equivalent to step 1 of Algorithm 1, then the two algorithms are equivalent. Recall that $Y^{(i)}_{\ell}=\bm{A}^{(i)}\bm{X}^{\ell-1}+\bm{G}^{(i)}_{\ell}$ and $\bm{G}^{(i)}_{\ell}\sim\mathcal{N}(0,\Delta_{l}^{2}\nu^{2})^{n\times p}$. Then steps 2 and 3 of Algorithm 2 correspond to:
$\begin{split}\bm{Y}_{\ell}&=SecAgg({\bm{Y}_{\ell}^{(i)}},\{i|1\leq i\leq s\})\\ &=\sum_{i=1}^{s}\bm{Y}_{\ell}^{(i)}\\ &=\sum_{i=1}^{s}(\bm{A}^{(i)}\bm{X}_{\ell-1}+\bm{G}^{(i)}_{\ell})\\ &=\bm{A}\bm{X}^{\ell-1}+\sum_{i=1}^{s}\bm{G}^{(i)}_{\ell}\\ &=\bm{A}\bm{X}^{\ell-1}+\bm{G}_{\ell},\end{split}$ (57)
where $\bm{G}_{\ell}\sim\mathcal{N}(0,\Delta_{l}^{2}\cdot(s\nu^{2}))^{n\times p}$, and we have $s\nu^{2}=\sigma^{2}$ by definition, completing the equivalence proof. ∎