### v2

# Differentially private and decentralized randomized power method

Abstract
The randomized power method has gained significant interest due to its simplicity and efficient handling of large-scale spectral analysis and recommendation tasks. However, its application to large datasets containing personal information (e.g. web interactions, search history, personal tastes) raises critical privacy concerns. This paper addresses these issues by proposing enhanced privacy-preserving variants of the method, which reduce the variance of the noise introduced to achieve Differential Privacy (DP). We also adapt the method to a decentralized framework with a low computational and communication overhead, while preserving the accuracy. We leverage Secure Aggregation to run the algorithm on data distributed among multiple users or devices, while using a noise scale in the decentralized setting that is similar to the one in the centralized setting yet without revealing individual data. We improve upon existing convergence bounds for both the centralized and decentralized versions.
Machine Learning, ICML
1 Introduction
The randomized power method is a simple and computationally efficient tool for solving large-scale linear algebra problems, particularly in spectral analysis and recommendation systems. It approximates the principal singular vectors of a matrix with low computational complexity (nearly linear), making it particularly suitable for analyzing large modern datasets.
In large-scale machine learning systems, protecting user privacy is paramount. Datasets often contain sensitive personal or organizational information, so privacy has become a critical requirement. The standard randomized power method does not inherently provide privacy guarantees. While its output might seem less sensitive than the input data, there is no formal guarantee that private information embedded in the data cannot be inferred.
To address and quantify privacy leakages, Differential Privacy (DP) has emerged as a powerful mathematical framework that provides strong formal guarantees and mitigate potential privacy leaks of an algorithm. DP ensures that the output of an algorithm reveals little about any individual record in the input. Several works have attempted to apply DP to the randomized power method. For instance, Hardt & Price (2014) and Balcan et al. (2016) introduced differentially private versions of the method, and Guo et al. (2021); Wang & Xu (2020) explored federated approaches to private power iteration. Other close works explore centralized differentially private versions of PCA (Liu et al., 2022) and federated (Wang & Xu, 2020; Briguglio et al., 2023) and give optimal convergence bounds given certain assumptions.
Despite these advancements, existing approaches suffer from several limitations. First, their performance heavily depends on the number of singular vectors being computed (Hardt & Price, 2014; Balcan et al., 2016; Guo et al., 2021; Liu et al., 2022; Wang & Xu, 2020), which impacts both utility and privacy guarantees. Second, they are primarily designed for centralized settings (Hardt & Price, 2014; Liu et al., 2022), where a trusted curator is assumed to hold the data. Moreover, some methods (Liu et al., 2022) make strong assumptions about the underlying data distribution (i.e. that the data is sub-Gaussian) which makes it harder to use these methods in practice. Some federated versions (Briguglio et al., 2023) claim to guarantee privacy by use of the federated setting, although it has been shown that decentralization does not offer privacy by design (Geiping et al., 2020). Finally, no fully decentralized versions exist to our knowledge. The previously proposed federated versions (Balcan et al., 2016; Guo et al., 2021; Wang & Xu, 2020) use public channel communication and ignore advances in Multi-Party-Computation (MPC). This makes them unsuitable for decentralized environments, such as distributed recommender systems and social networks, where data is partitioned across users and/or devices and communications are restricted to a predefined communication graph.
1.1 Contributions
The core focus of this article is to introduce improved privacy-preserving randomized power methods and to extend them to a decentralized setting, while maintaining computational efficiency and performance. We propose new convergence bounds for the (decentralized) privacy-preserving power method; these show that our proposed algorithm has a weaker dependence on the number of singular vectors that are computed.
We propose an improved differentially private version of the randomized power method, with tighter convergence bounds, reducing the noise required to achieve privacy and therefore improving the accuracy of the method. In the process of deriving these bounds, we present a new, more straightforward proof of a privacy result for the differentially private randomized power method. [^1] This proof addresses some minor mistakes in a privacy proof from Hardt & Roth (2012, 2013) that have been reproduced in several follow-up works (Hardt & Price, 2014; Balcan et al., 2016). The modified result allows us to use a wider range of privacy parameters.
Finally, we propose decentralized versions of the algorithms. These are particularly relevant in contexts where data is distributed across multiple users or devices, as is often the case in recommender systems. To address the privacy concerns of the users, without forcing them to rely on a trusted curator, we propose a federated privacy-preserving version of the method using Secure Aggregation (Bell et al., 2020), a form of Multi-Party Computation (MPC) and a fully decentralized version using GOPA (GOssip noise for Private Averaging) (Sabater et al., 2022). This allows users to keep their data local while contributing to a privacy-preserving global computation. Our decentralized approach preserves the effectiveness of the centralized version of the randomized power method while incorporating the privacy advantages of local differential privacy (DP). Notably, by employing Secure Aggregation, we are able to set the noise scale comparable to that of central DP, enhancing accuracy. This makes the proposed method a promising candidate for decentralized and privacy-conscious applications.
2 Background Material
2.1 Notation:
Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a real-valued positive semi-definite matrix, and let $k$ denote the target rank, where typically $k\ll n$. We also define $\eta$ as a small positive scalar that represents the precision of the approximation.
The eigenvalue decomposition of $\bm{A}$ is given by:
$\bm{A}=\bm{U}\bm{\Lambda}\bm{U}^{\top},$ (1)
where $\bm{U}$ is the matrix of eigenvectors of $\bm{A}$, and $\bm{\Lambda}$ is a diagonal matrix containing the corresponding eigenvalues on its diagonal.
Let $\mathcal{N}(\mu,\sigma^{2})^{n\times p}$ be a ($n\times p$) random matrix of which each element is a random variable independent and identically distributed (i.i.d.) according to a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$.
For any matrix $\bm{X}$, let its element-wise maximum norm be $\|\bm{X}\|_{\max}=\max_{i,j}\|\bm{X}_{ij}\|$, where $\bm{X}_{ij}$ is the $(i,j)$-th entry of matrix $\bm{X}$. We denote by $\bm{X}_{j:}$ the $j$-th row of $\bm{X}$.
We also define two coherence measures for the matrix $\bm{A}$:
  • The $\mu_{0}$-coherence of $\bm{A}$ is the maximum absolute value of its eigenvector entries, defined as $\mu_{0}(\bm{A})=\|\bm{U}\|_{\max}=\max_{i,j}|\bm{U}_{ij}|$.
  • The $\mu_{1}$-coherence of $\bm{A}$ is the maximum row sum of the absolute values of its eigenvector entries, defined as $\mu_{1}(\bm{A})=\|\bm{U}\|_{\infty}=\max_{i}(\sum_{j}|\bm{U}_{ij}|)$.
Finally, we will also use the QR decomposition of matrix obtained using the Gram-Schmidt procedure. Given a matrix $\bm{X}\in\mathbb{R}^{n\times p}$, the QR decomposition factorizes it as $\bm{X}=\bm{Q}\bm{R}$, where $\bm{Q}$ is an orthonormal matrix (i.e., $\bm{Q}^{\top}\bm{Q}=\bm{I}$) and $\bm{R}$ is an upper triangular matrix.
2.2 Differential privacy:
2.2.1 Approximate differential privacy:
A randomized algorithm $\mathcal{M}$ provides ($\epsilon,\delta$)-differential privacy if for all datasets $D_{1}$ and $D_{2}$ differing on at most one element (adjacent datasets), and for all measurable subsets $S\subseteq\text{Range}(\mathcal{M})$, the following holds:
$\Pr[\mathcal{M}(D_{1})\in S]\leq e^{\epsilon}\Pr[\mathcal{M}(D_{2})\in S]+\delta,$ (2)
where $\epsilon$ is a small positive parameter representing the privacy loss (smaller values indicate stronger privacy guarantees), and $\delta$ is a small probability representing the chance that the privacy guarantee may fail, allowing for a small probability of privacy breach.
2.2.2 $\ell_{2}$-sensitivity:
The $\ell_{2}$-sensitivity of a function $f$, denoted as $\Delta_{2}(f)$ or $\Delta_{2}$ for brevity, measures the maximum change in the output of $f$ when its input changes by at most one element (for neighboring datasets $D_{1}\sim D_{2}$). It is defined as:
$\Delta_{2}=\max_{D_{1}\sim D_{2}}\|f(D_{1})-f(D_{2})\|_{2},$ (3)
where $\|\cdot\|_{2}$ represents the $\ell_{2}$-norm.
2.2.3 Zero-Concentrated Differential Privacy (zCDP):
A randomized algorithm $\mathcal{M}$ is said to satisfy $\rho$-zero-Concentrated Differential Privacy (zCDP) if for all neighboring datasets $D_{1}$ and $D_{2}$, and for all $\alpha\in(1,\infty)$, the following holds:
$D_{\alpha}(\mathcal{M}(D_{1})\|\mathcal{M}(D_{2}))\leq\rho\alpha\,,$ (4)
where $D_{\alpha}$ is the Rényi divergence of order $\alpha$ and $\rho$ is a positive parameter controlling the trade-off between privacy and accuracy (smaller values of $\rho$ imply stronger privacy guarantees).
The following lemma, introduced by Bun & Steinke (2016), specifies how the addition of Gaussian noise can be used to construct a randomized algorithm that satisfies zCDP.
Lemma 2.1 (Gaussian Mechanism (Proposition 1.6 (Bun & Steinke, 2016))).
Let $f:X^{n}\to\mathbb{R}$ be a sensitivity-$\Delta$ function. Consider the mechanism (randomized algorithm) $M:X^{n}\to\mathbb{R}$, defined as $M(x)=f(x)+Z_{x}$ where for each input $x$, $Z_{x}$ is independently drawn from $\mathcal{N}(0,\sigma^{2})$. Then $M$ satisfies $\left(\frac{\Delta^{2}}{2\sigma^{2}}\right)$-zCDP.
The next lemma, which is a generalized version of a result presented by Bun & Steinke (2016), explains how a randomized algorithm, constructed by recursively composing a sequence of zCDP-satisying randomized algorithms, also satisfies zCDP.
Lemma 2.2 (Adaptive composition (Generalization from Lemma 2.3 (Bun & Steinke, 2016))).
Let $M_{1}:X^{n}\to Y_{1}$, $M_{2}:X^{n}\times Y_{1}\to Y_{2}$, …, $M_{L}:X^{n}\times Y_{1}\times\dots\times Y_{L-1}\to Y_{L}$ be randomized algorithms. Suppose $M_{i}$ satisfies $\rho_{i}$-zCDP as a function of its first argument for each $i=1,2,\dots,L$. Let $M^{\prime\prime}:X^{n}\to Y_{L}$, constructed recursively by:
$M^{\prime\prime}(x)=M_{L}(x,M_{L-1}(x,\dots,M_{2}(x,M_{1}(x))\dots)).$ (5)
Then $M^{\prime\prime}$ satisfies $(\sum_{i=1}^{L}\rho_{i})$-zCDP.
2.3 Privacy preserving power iteration:
Let $k$ denote the target rank, $q$ a small real and $\eta$ a small positive value controlling the precision of the approximation. The goal of the randomized power method is to compute a matrix $\bm{Q}\in\mathbb{R}^{n\times p}$, where $p=k+q$, such that the range of $\bm{Q}$ approximates the range of the top $k$ eigenvectors of $\bm{A}$, which are denoted $\bm{U}_{k}$. Specifically, we aim to ensure that:
$\|\bm{U}_{k}-\bm{Q}\bm{Q}^{\top}\bm{U}_{k}\|\leq\eta.$
The privacy-preserving randomized power method extends this approach to protect sensitive information in $\bm{A}$ during computation. Prior work (Hardt & Price, 2014; Balcan et al., 2016) show that under approximate Differential Privacy (DP) parameters $(\epsilon,\delta)$, $\bm{Q}$ can be computed using Algorithm 1, where the sensitivity $\Delta_{l}$ at iteration $l$ is bounded by $\Delta_{l}^{\text{prior}}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$. Here, $\Delta_{l}^{\text{prior}}$ represents an $\ell_{2}$-sensitivity bound that quantifies the impact of a single-element change in $\bm{A}$ on $\bm{A}\bm{X}^{l-1}$. Specifically, $\Delta_{l}\leq\Delta_{l}^{\text{prior}}$.
With this bound [^2] As defined in Fig. 3 of (Hardt & Price, 2014), Alg. 2 of (Balcan et al., 2016) and Theorem 6 of (Guo et al., 2021)., the standard deviation of the noise $\sigma_{l}$ associated to DP at iteration $l$ is given by:
$\sigma_{l}$ $=\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$ (6)
$=\Delta_{l}^{\text{prior}}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}.$ (7)
As a first contribution, we refine this sensitivity analysis by introducing a significantly tighter bound for $\Delta_{l}$, reducing the noise required for DP and therefore allowing tighter convergence bounds (smaller $\eta$).
1:  Input: Matrix $\bm{A}\in\mathbb{R}^{n\times n}$, number of iterations $L$, target rank $k$, iteration rank $p\geq k$, privacy parameters $\epsilon,\delta$.
2:  Output: approximated eigen-space $\bm{X}^{L}\in\mathbb{R}^{n\times p}$, with orthonormal columns.
3:  Initialization: orthonormal $\bm{X}^{0}\in\mathbb{R}^{n\times p}$ by QR decomposition on a random Gaussian matrix $\bm{G}_{0}$; noise variance parameter $\sigma=\epsilon^{-1}\sqrt{{4L\log(1/\delta)}}$;
4:  for $\ell=1$ to $L$ do
5:      1. Compute $\bm{Y}_{\ell}=\bm{A}\bm{X}^{\ell-1}+\bm{G}_{\ell}$ with $\bm{G}_{\ell}\sim\mathcal{N}(0,\sigma_{l}^{2}={\Delta}_{l}^{2}\cdot\sigma^{2})^{n\times p}$2. Compute QR factorization $\bm{Y}_{\ell}=\bm{X}^{\ell}\bm{R}_{\ell}$.
6:  end for
Algorithm 1 Privacy-preserving randomized power method
2.4 Existing convergence bounds
To our knowledge, the strongest bound for the randomized power method as defined in Algorithm 1 with $\Delta_{l}=\Delta_{l}^{prior}\triangleq\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$ is given in Balcan et al. (2016). In Theorem 2.4, we present a reduction of their Theorem 3.1. The theorem, and several others we present relies on the verification of the conditions in Assumption 2.3.
Assumption 2.3.
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$, and iteration rank $p$, with $k\leq q\leq p$. Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues. Let us fix:
$\eta=O\left(\frac{\sigma_{q}}{\sigma_{k}}\cdot\min\left\{\frac{1}{\log\left(\frac{\sigma_{k}}{\sigma_{q}}\right)},\frac{1}{\log\left(\tau d\right)}\right\}\right).$ (8)
Assume that at every iteration $l$ of Algorithm 1, $\bm{G}_{\ell}$ satisfies, for some constant $\tau>0$:
$\|\bm{G}_{\ell}\|_{2}$ $=O\left(\eta(\sigma_{k}-\sigma_{q+1})\right)\,,\quad\text{and}$ (9)
$\quad\|\bm{U}_{q}^{\top}\bm{G}_{\ell}\|_{2}$ $=O\left(\eta\left(\sigma_{k}-\sigma_{q+1}\right)\frac{\sqrt{p}-\sqrt{q-1}}{\tau\sqrt{d}}\right)\,.$ (10)
Theorem 2.4 (Private Power Method (PPM) (Balcan et al., 2016)).
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$, and iteration rank $p$, with $2q\leq p\leq d$. Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$. Let $\epsilon,\delta\in(0,1)$ be the differential privacy parameters. Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues. Then if Assumption 2.3 holds, Algorithm 1 with $\Delta_{l}=\Delta_{l}^{prior}$ gives us the following runtime-dependent convergence bound:
$\eta=O\left(\frac{\epsilon^{-1}\max_{l}{(\|\bm{X}_{l}\|_{\max})}\sqrt{pLd\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right)$
and also the following runtime-independent convergence bound
$\eta=O\left(\frac{\epsilon^{-1}\|\bm{U}\|_{\max}\sqrt{pLd\log(1/\delta)\log(d)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$
3 Analysis of Runtime-dependent Bounds
3.1 Proposed runtime-dependent bounds
3.1.1 Adjacency notion:
In prior works (Hardt & Price, 2014; Balcan et al., 2016), two matrices $\bm{A}$ and $\bm{A}^{\prime}$ are considered adjacent (denoted $\bm{A}\sim\bm{A}^{\prime}$) if they differ by a change in a single element, with a Frobenius norm of at most 1:
$\bm{A}^{\prime}=\bm{A}+c\cdot\bm{e}_{i}\bm{e}_{j}^{\top},$ (11)
where $\bm{e}_{i},\bm{e}_{j}\in\mathbb{R}^{n}$ are canonical basis vectors of dimension $d$, $c\leq 1$ represents the magnitude of the change, and $0\leq i,j<n$. This adjacency notion models a sensitive change as a modification to a single element in $\bm{A}$ to protect individual element-wise updates under Differential Privacy. However, since $\bm{A}$ is symmetric positive semi-definite and changes must preserve this property, this necessitates updates on the diagonal elements. We propose a new, more general, notion of adjacency to allow for other types of updates:
$\bm{A}^{\prime}=\bm{A}+\bm{C},$ (12)
where $\bm{C}\in\mathbb{R}^{n\times n}$ is a symmetric matrix representing the update, subject to the condition:
$\sqrt{\sum_{i=1}^{n}\|\bm{C}_{i,:}\|_{1}^{2}}\leq 1.$ (13)
The proposed adjacency notion is strictly more general than Equation (11). For example, setting $\bm{C}=c\cdot\bm{e}_{i}\bm{e}_{j}^{\top}$ recovers the original definition. Additionally, $\bm{C}$ can have non-zero entries on the diagonal, anti-diagonal, or any symmetric pattern, allowing for a variety of updates while maintaining symmetry.
In the context of recommender systems where $\bm{A}=\bm{R}^{\top}\bm{R}$ and $\bm{R}$ is a user-item interaction matrix, this adjacency notion allows for element-wise changes in $\bm{R}$ (i.e. protect user-item interactions). By adopting this more general adjacency definition, we make our privacy guarantees applicable to a wider range of real-world scenarios.
3.1.2 Improved sensitivity bound:
The previous sensitivity bound $\Delta_{l}^{\text{prior}}$ (Hardt & Price, 2014; Balcan et al., 2016) for $\bm{A}\bm{X}^{l-1}$ was estimated by extrapolating from the case where $\bm{X}^{l-1}\in\mathbb{R}^{n\times 1}$ to the general case $\bm{X}^{l-1}\in\mathbb{R}^{n\times p}$. This leads to an overestimation of the sensitivity and to unnecessarily large noise addition. We directly analyze the change in $\bm{X}^{l-1}\in\mathbb{R}^{n\times p}$ and derive a tighter bound on the sensitivity $\Delta_{l}$. By using our proposed adjacency notion from Equation (12), we establish the following result (proof provided in Appendix A.1):
Theorem 3.1 (Improved Sensitivity Bound).
Let $\bm{A}^{\prime}$ be defined as in Equation (12), and consider the sensitivity $\Delta_{l}=\|\bm{A}^{\prime}\bm{X}^{l}-\bm{A}\bm{X}^{l}\|_{F}$. Then, $\Delta_{l}$ is upped-bounded as follows:
$\Delta_{l}\leq\max_{i}\|\bm{X}_{i,:}^{l}\|_{2}\triangleq\hat{\Delta}_{l}.$ (14)
This improved bound $\hat{\Delta}_{l}$ is generally tighter than the previous bound $\Delta_{l}^{\text{prior}}=\sqrt{p}\cdot\|\bm{X}_{l}\|_{\max}$, as it eliminates the dependence on $\sqrt{p}$ and directly relates to the maximum row norms of $\bm{X}^{l}$. By reducing the estimated sensitivity, we can add less noise while still satisfying Differential Privacy, thereby improving the utility of the algorithm.
3.1.3 Privacy proof:
To address these issues, we propose a result with a new proof of the Differential Privacy guarantees for our algorithm in Theorem 3.2, whose proof is in Appendix A.2.2. This proof clarifies the ambiguities or errors present in the previous works (see Appendix A.2), and ensures that our privacy guarantees are met.
Theorem 3.2 (Privacy proof for the PPM).
Assuming $\delta\in(0,1)$ and $\epsilon>0$ such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$, the Private Power Method with $\Delta_{l}=\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ satisfies ($\epsilon,\delta$)-Differential Privacy.
3.1.4 Improved convergence bound:
Theorem 3.3 (Bound for the noisy power method (NPM) (Balcan et al., 2016)).
Let reals $k\leq q\leq p$. Let $\bm{U}_{q}\in\mathbb{R}^{d\times q}$ be the top-$q$ eigenvectors of a positive semi-definite matrix $\bm{A}$ and let $\sigma_{1}\geq\cdots\geq\sigma_{d}\geq 0$ denote its eigenvalues and fix $\eta=O\left(\frac{\sigma_{q}}{\sigma_{k}}\cdot\min\left\{\frac{1}{\log\left(\frac{\sigma_{k}}{\sigma_{q}}\right)},\frac{1}{\log\left(\tau d\right)}\right\}\right)$. If at every iteration $l$ of the NPM $\bm{G}_{\ell}$ satisfies Assumption 2.3, then after
$L=\Theta\left(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log\left(\frac{\tau d}{\eta}\right)\right).$
iterations, with probability at least $1-\tau^{-\Omega(p+1-q)}-e^{-\Omega(d)}$, we have:
$\|(\bm{I}-\bm{X}^{L}{\bm{X}^{L}}^{\top})\bm{U}_{k}\|_{2}\leq\eta.$ (15)
Theorem 3.4 (Improved PPM with runtime dependent bound).
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$ and iteration rank $p$ with $2q\leq p\leq d$. Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$. Let $\delta\in(0,1)$ and $\epsilon>0$ be privacy parameters such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$
Then with approximate DP parameters ($\epsilon$,$\delta$) and assuming Assumption 2.3 is verified, Algorithm 1 with $\Delta_{l}=\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ gives us the following runtime-dependent convergence bound:
$\eta=O\left(\frac{\epsilon^{-1}\max_{i,l}\|\bm{X}_{i:}^{l}\|_{2}\sqrt{Ld\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$ (16)
Note: This bound is at least as tight as the one proposed in (Balcan et al., 2016), and can be in practice much tighter depending on $\bm{A}$. Our privacy proof (given in Theorem 3.2) does not restrict us to have $\epsilon\leq 1$. We provide the proof in Appendix A.2.2.
4 Analysis of Runtime-independent bounds
4.1 Proposed runtime-independent convergence bound
We presented in Theorem 3.4 a runtime-dependent convergence bound involving $\max_{i}\|\bm{X}_{i,:}\|_{2}$, which is only known during the execution of the algorithm. To provide a more practical analysis, we now derive a runtime-independent convergence bound in Theorem 4.1. This bound is a combination of two cases:
  • High Coherence Case: Applies to matrices with large coherence parameters, and the bound is independent of the matrix’s coherence.
  • Low Coherence Case: Applies to matrices with low $\mu_{1}$-coherence, with the bound depending explicitly on $\mu_{1}(\bm{A})$.
We demonstrate that our method achieves a convergence bound that is at least as tight as both of these bounds.
We note that for a matrix $\bm{A}$, the proposed bound is tighter than the one proposed in (Balcan et al., 2016) when $\min(1,\mu_{1}(\bm{A}))\leq\sqrt{p\cdot\log({d})}\cdot\mu_{0}(\bm{A})$. This is likely to happen in practice, as we highlight in Section 6.1.1.
Theorem 4.1.
Improved PPM with runtime independent bound
Let $\bm{A}\in\mathbb{R}^{d\times d}$ be a symmetric data matrix. Fix target rank $k$, intermediate rank $q\geq k$ and iteration rank $p$ with $2q\leq p\leq d$. Suppose the number of iterations $L$ is set as $L=\Theta(\frac{\sigma_{k}}{\sigma_{k}-\sigma_{q+1}}\log(d))$. Let $\delta\in(0,1)$ and $\epsilon>0$ be privacy parameters such that $\delta\leq\exp{(-\frac{\epsilon}{4})}$ Then with approximate DP parameters ($\epsilon$,$\delta$) and assuming Assumption 2.3 is verified, Algorithm 1 gives us the following runtime-independent convergence bound:
$\eta=O\left(\frac{\epsilon^{-1}\cdot\min(1,\mu_{1}(\bm{A}))\cdot\sqrt{Ld\log(1/\delta)\log(L)}}{\sigma_{k}-\sigma_{q+1}}\right).$ (17)
5 Decentralized version
Table 1: Overhead of using Secure Aggregation in our protocol
Client Server
Communication $O(log(s)+n^{2})$ $O(s\cdot log(s)+sn^{2})$
Computation $O(log^{2}(s)+n^{2}\cdot log(s))$ $O(s\cdot log^{2}(s)+sn^{2}\cdot log(s))$
In this section, we consider a decentralized setting where the matrix $\bm{A}$ is distributed across multiple clients. Specifically, each client $i$ holds a private matrix $\bm{A}^{(i)}\in\mathbb{R}^{n\times n}$, such that the global matrix is the sum of these local matrices: $\bm{A}=\sum_{i=1}^{s}\bm{A}^{(i)}.$
The goal is for the clients to collaboratively compute an orthonormal basis for the range of $\bm{A}$, similar to the centralized randomized power method, but without revealing their individual private matrices $\bm{A}^{(i)}$ to the server or to other clients.
The randomized power method involves linear operations, making it well-suited for parallelization and distributed computation. Balcan et al. (2016) proposed a private and federated power method using communication over public channels between clients and a server. However, this approach relies on Differential Privacy, since the data exchanged can be observed by everyone, which requires high levels of noise to ensure privacy.
To enhance privacy while retaining the benefits of distributed computation, we use Secure Aggregation, a lightweight Multi-Party Computation protocol. Secure Aggregation allows clients to collaboratively compute sums without revealing individual data, enabling the use of distributed DP. This approach offers privacy guarantees similar to central DP and eliminates the need for a trusted curator. Distributed DP has been extensively studied in the literature (Chen et al., 2021; Kairouz et al., 2021; Goryczka et al., 2013; Wei et al., 2024; Ghazi et al., 2019).
In distributed DP, each client adds carefully calibrated noise to their local contributions before participating in the Secure Aggregation protocol. The noise is designed such that the sum across clients has a variance comparable to that used in central DP, thereby achieving similar Differential Privacy guarantees without requiring a trusted aggregator, and allowing clients to keep the data locally.
We operate under the “honest-but-curious“ threat model, where clients follow the protocol correctly but may attempt to learn additional information from received data. We also assume there are no corrupted or dropout users during the computation. For simplicity, this paper focuses on a basic implementation of distributed DP, neglecting the effects of data quantization or errors introduced by modular arithmetic. We refer to Kairouz et al. (2021) for an implementation of Distributed DP that takes these effects into account.
Finally, it is possible to use a fully decentralized and differentially private protocol to perform averaging without relying on a central server for Secure Aggregation, as demonstrated by Sabater et al. (2022). Adopting such approaches can further enhance decentralization while maintaining similar utility and communication costs.
5.1 Decentralized and private power method using distributed DP:
We introduce a federated version of the Privacy-Preserving Power Method in Algorithm 2. This version reduces the noise variance by a factor of $O(sp\log d\,\|\bm{U}\|_{\infty}^{2})$ compared to the method in Balcan et al. (2016). The improvement comes from the fact that we emulate centralized Differential Privacy (DP) over secure channels, avoiding the higher noise required in local DP settings with public channels, and that we use our generally tighter sensitivity bound.
We demonstrate the equivalence between the decentralized Algorithm 2 and the centralized Algorithm 1 in Theorem 5.2 (proof provided in Appendix A.5). We also give a simplified definition of the Secure Aggregation algorithm.
Definition 5.1.
We denote by $SecAgg({\bm{Y}_{\ell}^{(i)}},\{i|1{\leq}i{\leq}s\})$ the Secure Aggregation of locally held matrices $\bm{Y}_{\ell}^{(i)}$ over the set of users represented by indices $\{i|1{\leq}i{\leq}s\})$. It is equivalent to computing $\bm{Y}_{\ell}=\sum_{i=1}^{s}\bm{Y}_{\ell}^{(i)}$ over a secure channel.
  Input: distributed matrices $\bm{A}^{(1)},\cdots,\bm{A}^{(s)}\in\mathbb{R}^{n\times n}$, number of iterations $L$, target rank $k$, iteration rank $p\geq k$, private parameters $\epsilon,\delta$.
  Output: approximated eigen-space $\bm{X}^{L}\in\mathbb{R}^{n\times p}$, with orthonormal columns.
  Initialization: orthonormal $\bm{X}^{0}\in\mathbb{R}^{n\times p}$ by QR decomposition on a random Gaussian matrix $\bm{G}_{0}$; noise variance parameter $\nu=\epsilon^{-1}\sqrt{\frac{4L\log(1/\delta)}{s}}$;
  for $\ell=1$ to $L$ do
      1. The central node broadcasts $\bm{X}^{\ell-1}$ to all $s$ computing nodes;2. Computing node $i$ computes $\bm{Y}^{(i)}_{\ell}=\bm{A}^{(i)}\bm{X}^{\ell-1}+\bm{G}^{(i)}_{\ell}$ with $\bm{G}^{(i)}_{\ell}\sim\mathcal{N}(0,\Delta_{l}^{2}\nu^{2})^{n\times p}$;3. The central node computes with the clients $\bm{Y}_{\ell}=SecAgg({\bm{Y}_{\ell}^{(i)}},\{i|1\leq i\leq s\})$. 4. The central node computes QR factorization $\bm{Y}_{\ell}=\bm{X}^{\ell}\bm{R}_{\ell}$.
  end for
Algorithm 2 Federated private power method
Theorem 5.2 (Privacy and utility of Algorithm 2).
The decentralized Privacy-Preserving Power Method (Algorithm 2) provides the same privacy guarantees and achieves equivalent utility (in terms of convergence) as its centralized version (Algorithm 1).
5.2 Communication and computation cost analysis:
We denote by $s$ the number of shards, such that $\sum_{i=1}^{s}\bm{A}^{(i)}=\bm{A}\in\mathbb{R}^{n\times n}$. Table 1 presents the communication and computations cost overhead introduced by the polylogarithmic version of Secure Aggregation that we employ (Bell et al., 2020).
6 Empirical comparison of the proposed bounds
We introduced both runtime-dependent and runtime-independent convergence bounds for our algorithm. The runtime-dependent bound depends on the matrix $\bm{A}$ and the iteratively computed matrices $\bm{X}_{l}$, whereas the runtime-independent bound (Theorem 4.1) only depends on $\bm{A}$. To illustrate the practical impact of our changes, we focus on an application in recommender systems and compare how our algorithm perform in this context. We will then use a statistical approximation to see how the runtime-dependent bounds compare at the first step of the algorithm, regardless of the application.
6.1 Application to recommender systems
State-of-the-art recommender systems like GF-CF (Shen et al., 2021) and BSPM (Choi et al., 2023) utilize singular value decomposition (SVD) or eigenvalue decomposition as part of their algorithms. Specifically, they represent the dataset (comprised of user-item interactions) using a normalized adjacency matrix $\tilde{\bm{R}}$, from which they compute a normalized item-item matrix $\tilde{\bm{P}}$. To get rid of noise in $\tilde{\bm{R}}$, they compute an ideal low-pass filter based on the top-$p$ eigenvectors of $\tilde{\bm{P}}$ and apply it to $\tilde{\bm{R}}$. Detailed definitions are provided in Appendix A.7.
Table 2 presents key statistics of popular recommendation datasets used in the literature. Depending on the use case, these datasets may be held by a central curator, or distributed among the users of a recommender system. Accordingly, either the centralized PPM (Algorithm 1) or the decentralized PPM (Algorithm 2) can be applied to compute the leading eigenvectors. To demonstrate the practicality of our proposed methods, we focus on a decentralized setting where each user has access to part of the dataset and collaborate with other users to compute the desired eigenvectors.
Table 2: Statistics of datasets
Dataset Users Items Interactions $\mu_{0}$ $\mu_{1}$ $\Delta_{l}^{prior}$ (existing) ${\hat{\Delta}_{l}}$ (ours)
Amazon-book 52,643 91,599 2,984,108 0.33 52.95 1.12 ×$\times\sqrt{p}$ 1
MovieLens 71,567 10,677 7,972,582 0.39 32.84 1.19 ×$\times\sqrt{p}$ 1
EachMovie 74,425 1,649 2,216,887 0.49 21.89 1.33 ×$\times\sqrt{p}$ 1
Jester 54,906 151 1,450,010 0.73 7.78 1.64 ×$\times\sqrt{p}$ 1
6.1.1 Empirical comparison of the runtime-independent bounds
As discussed in Section 4.1, our proposed runtime-independent bound is tighter than those proposed in Hardt & Price (2014); Balcan et al. (2016) when $\min(1,\mu_{1}(\bm{A}))\leq\mu_{0}(\bm{A})\sqrt{p\cdot\log({d})}$, where $\bm{U}$ denotes the eigenvectors of $\bm{A}$. Table 2 shows that this condition holds for popular recommender system datasets. As we can deduce from Table 2, the proposed bound theoretically allows us to converge to solutions with much smaller $\eta$ values, especially when the desired number of factors $p$ is large.
6.1.2 Practical utility of the proposed algorithms
We saw that the proposed runtime-independent bounds were practically tighter than the previous one for the task of interest. We demonstrate that our decentralized PPM can privately compute the top-$p$ eigenvectors $\bm{U}_{p}$ of the user-normalized item-item matrix $\tilde{\bm{P}}$ (proof in Appendix A.1):
Lemma 6.1.
Algorithm 2 with $\Delta_{l}=\sqrt{2}\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ can compute $\bm{U}_{p}$, the top-$p$ eigenvectors of $\tilde{\bm{P}}$ in a decentralized setting under $(\epsilon,\delta)$-Differential Privacy.
Figure 1: Impact of the Differential Privacy parameter $\epsilon$ on the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to multiple PPM versions, computed with $p=32$ and $L=3$ on the EachMovie dataset.
Both GF-CF (Shen et al., 2021) and BSPM (Choi et al., 2023) use $\bm{U}_{p}$ to compute the ideal low pass filter and apply it to the interaction matrix $\bm{R}$, giving a filtered interaction matrix $\bm{R}_{p}$. To illustrate the practicality of our proposed method, we compute approximations of $\bm{U}_{p}$ (denoted by $\tilde{\bm{U}}_{p}$) using either our proposed decentralized PPM, or the PPM versions proposed by  Balcan et al. (2016). We then use $\tilde{\bm{U}}_{p}$ to compute an approximation of $\bm{R}_{p}$ (noted $\tilde{\bm{R}_{p}}$), and compute the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to each version.
Figure 1 illustrates the impact of the privacy parameter $\epsilon$ on the approximation error for the EachMovie dataset, with $p=32$ and $L=3$. Our decentralized PPM achieves a relatively low approximation error (approximately 1/10 of the random guessing error) for practical values of $\epsilon$ (5-10). In contrast, the methods from Balcan et al. (2016) require $\epsilon$ values of at least $10^{3}$ to achieve comparable errors, which does not provide meaningful privacy protection. Additional comparisons and details are provided in Appendix A.7.1.
6.2 Matrix-agnostic comparison of the runtime-dependent bounds
We analyze the tightness of our proposed runtime-dependent bound compared to the previous bound in a matrix-agnostic manner, focusing on the first iteration of the algorithm.
At iteration $l=0$, the leading singular vectors are initialized as $\bm{X}^{0}=\text{Q}(\bm{\Omega})$, where $\bm{\Omega}\sim\mathcal{N}(0,1)^{n\times k}$ and $\text{Q}(\bm{\Omega})$ denotes the orthonormal matrix obtained from the QR decomposition of $\bm{\Omega}$. Since $\bm{X}^{0}$ is independent of the matrix $\bm{A}$, it allows us to assess the relative tightness of the proposed bound $\hat{\Delta}_{0}$ compared to the previous bound $\Delta_{0}^{prior}$ at the first step.
Let $r(k,n)$ be the ratio of the expected values of the two bounds, as a function of the desired number of eigenvectors $k$ and the matrix dimension $n$:
$r(k,n)=\frac{\mathbb{E}[\Delta_{0}^{prior}]}{\mathbb{E}[\hat{\Delta}_{0}]}.$ (18)
6.2.1 Theoretical approximation of $r(k,n)$:
We derive an asymptotic approximation of $r(k,n)$ using Approximation 6.2 (detailed in Appendix A.4).
Approximation 6.2.
[…]
$\mathbb{E}[\Delta_{l}^{prior}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(kn)\frac{k^{2}}{n^{2}}}+\frac{k\mu}{n}\right)$ (19)
$\mathbb{E}[\hat{\Delta}_{l}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(n)\frac{k}{n^{2}}}+\frac{k\mu}{n}\right).$ (20)
6.2.2 Empirical approximation of $r(k,n)$:
Figure 2: Comparison of empirically (in blue) and theoretically (in green) estimated $r(k,n)$ ratios, with $n=8000$ and $k$ ranging between 64 and 4000, with a step of 64. Blue dots represent empirical ratios estimated using $t=5$ runs of the first step of our algorithm. The theoretical ratios are estimated using Theorem 6.2.
To compare the tightness of our proposed bound $\hat{\Delta}_{0}$ to the previous bound $\Delta_{0}^{prior}$, we perform both theoretical and empirical analyses at the first iteration, where the noise scaling depends only on the random initialization $\bm{X}^{0}$. Since $\bm{X}^{0}$ is independent of $\bm{A}$, we can estimate $\mathbb{E}[\Delta_{0}^{prior}]$ and $\mathbb{E}[\hat{\Delta}_{0}]$ by sampling random Gaussian matrices $\bm{X}^{0}$.
Experiment: $r(k,n)$ depends on the number of factors $k$ and on $n$, where ($n\times n$) is the size of $\bm{A}$. We therefore seek to compute it for multiple values of $n$ ($n\in\{8000,12000\}$) and a range of values for $k$ (between 64 and 4000, with steps of 64), for each value of $n$. We use Theorem 6.2 to approximate it theoretically. To measure it empirically, we compute $\Delta_{0}^{prior}$ and $\hat{\Delta}_{0}$ based on $\bm{X}_{0}$ in the first step of the algorithm (it is independent of the matrix of interest $\bm{A}$), for $t$ runs of the algorithm. We can then use the $t$ measures to estimate $\mathbb{E}[\Delta_{0}^{prior}]$ and $\mathbb{E}[\hat{\Delta}_{0}]$. We claim that $t=5$ is enough to see a general tendancy, as we have estimates for many ($k,n$) couples. We can also compare the empirical estimates to the proposed asymptotic approximations specified in Approximation 6.2.
Figure 2 and Figure 3 (in Appendix) presents the comparison between empirical (blue dots) and theoretical (green line) estimates of $r(k,n)$ for $n=8000$ and $n=12000$. The empirical ratios are calculated from the averages over the $t=5$ runs for each value of $k$. The results indicate that both empirically and theoretically the proposed noise scaling $\hat{\Delta}_{l}$ is much tighter than $\Delta_{l}^{prior}$ at the first step of the algorithm. Our theoretical approximation is conservative and underestimates how much tighter the bound is initially, compared to what we observe in practice. The proposed bound is tighter by a multiplicative factor on the first step and can therefore drastically reduce the impact of the noise introduced by DP on the convergence of the Privacy Preserving power method, as the algorithm is usually run for very few steps ($L$ is usually in the range of $1$-$5$).
7 Conclusion and limitations
We presented a differentially private and decentralized randomized power method that addresses privacy concerns in large-scale spectral analysis and recommendation systems. We introduced new convergence bounds, improving the accuracy of the method while ensuring privacy guarantees. By employing Secure Aggregation in a decentralized setting, we were able to reduce the noise introduced for privacy, maintaining the efficiency and privacy of the centralized version but adapting it for distributed environments.
However, our privacy guarantees rely on the honest-but-curious threat model and assume no collusion or malicious behavior from participants, which may not always be realistic. In the future, it would also be interesting to analyze a private and decentralized version of the accelerated version of the randomized power method, which could potentially have a faster convergence.
8 Impact Statement (required by ICML)
Our contributions directly address the need for privacy-preserving algorithms in the context of large-scale machine learning systems, where sensitive personal or organizational data is often at stake. The decentralized implementation ensures that the method can be applied in some real-world environments where data is distributed across multiple users or devices, making it particularly suitable for recommender systems and other distributed applications. This work is expected to have implications for privacy-conscious machine learning in real-world systems, to perform large-scale data analysis while maintaining robust privacy protections.
Appendix A Proofs of Results
A.1 Proof of Theorem 3.1
Proof.
We note $\bm{A}^{\prime}$ a matrix adjacent to $\bm{A}$ using (12).
$\bm{A}^{\prime}=\bm{A}+\bm{C},$ (21)
with $\bm{C}$ a symmetric matrix representing the update, subject to $\sqrt{\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}}\leq 1$.
Then
$\begin{split}\Delta_{l}&=\|\bm{A}\bm{X}^{l}-\bm{A}^{\prime}\bm{X}^{l}\|_{F}\\ &=\|\bm{A}\bm{X}^{l}-(\bm{A}+\bm{C})\bm{X}^{l}\|_{F}\\ &=\|\bm{C}\bm{X}^{l}\|_{F}\\ &=\sqrt{\sum_{i}\|\bm{C}_{i:}\bm{X}^{l}\|_{F}^{2}}.\\ \end{split}$ (22)
and
$\begin{split}\|\bm{C}_{i:}\bm{X}^{l}\|_{F}&=\|\sum_{j}\bm{C}_{ij}\cdot\bm{e}_{j}^{\top}\cdot\bm{X}^{l}\|_{F}\\ &=\|\sum_{j}\bm{C}_{ij}\cdot\bm{X}^{l}_{j:}\|_{F}\\ &\leq\sum_{j}\|\bm{C}_{ij}\cdot\bm{X}^{l}_{j:}\|_{F}\\ &\leq\sum_{j}|\bm{C}_{ij}|\cdot\|\bm{X}^{l}_{j:}\|_{F}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\cdot\sum_{j}|\bm{C}_{ij}|.\end{split}$ (23)
By injecting (22) into (23), we have:
$\begin{split}\Delta_{l}&=\sqrt{\sum_{i}\|\bm{C}_{i:}\bm{X}^{l}\|_{F}^{2}}\\ &\leq\sqrt{\sum_{i}(\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\cdot\sum_{j}|\bm{C}_{ij}|)^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}(\cdot\sum_{j}|\bm{C}_{ij}|)^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\triangleq\hat{\Delta}_{l}.\\ \end{split}$ (24)
Although the proposed update model (12) is more general than when $\bm{A}^{\prime}$ is defined using (11), the proposed bound $\hat{\Delta}_{l}$ is also generally tighter than the bound proposed in (Hardt & Price, 2014; Balcan et al., 2016).
A.2 Privacy proof
A.2.1 Note on related privacy proofs:
Several Differential Privacy (DP) proofs for the private randomized power method (PPM) have been developed in prior works, for instance those by Hardt & Price (2014), Balcan et al. (2016), Hardt & Roth (2013), Hardt & Roth (2012). These papers mostly rely on the same proofs to establish that the PPM satisfies ($\epsilon$, $\delta$)-DP. Specifically, Balcan et al. (2016) references the privacy proof from Hardt & Price (2014), which builds upon the privacy proof from Hardt & Roth (2013). The proof therein is also closely related to the one of Hardt & Roth (2012).
However, both Hardt & Roth (2013) (Theorem 2.4) and Hardt & Roth (2012) (Theorem 2.4) contain errors in their proposed composition rules, where a comparison sign is mistakenly flipped. This error could potentially cause the privacy parameter $\epsilon$ in the proposed mechanism to be arbitrarily small, providing no privacy guarantee at all. The original composition rule is presented in Theorem III.3 of Dwork et al. (2010). Moreover, Lemma 3.4 of Hardt & Roth (2013) misuses Theorem 2.4. Indeed, they claim that their algorithm satisfies $(\epsilon^{\prime},\delta)$-DP at each iteration. By their Theorem 2.4, then the algorithm overall should satisfy $(\epsilon^{\prime},k\delta+\delta^{\prime})$-DP where $\delta^{\prime}>0$, and not $(\epsilon^{\prime},\delta)$-DP as claimed.
A.2.2 Proof of Theorem 3.2
Proof.
We can see lines 4-5 of Algorithm 1 as a sequential composition ($M$) of $L$ Gaussian Mechanisms. By Lemma 2.1, each mechanism $M_{i}$ satisfies ($\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$)-zCDP where $\Delta_{i}$ is the $\ell_{2}$-sensitivity of the function associated to mechanism $\Delta_{i}$ and $\sigma_{i}^{2}$ is the variance of the noise added with the Gaussian Mechanism. By Lemma 2.2, the composition of mechanisms $M=(M_{1},...,M_{i},...,M_{L})$ satisfies ($\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$)-zCDP. Let $\rho\triangleq\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$. By design of our algorithm, we have:
$\rho$ $=\sum_{i=1}^{L}\frac{\Delta_{i}^{2}}{2\sigma_{i}^{2}}$ (25)
$\leq\sum_{i=1}^{L}\frac{(\hat{\Delta}_{l})^{2}}{2\sigma_{i}^{2}}$ (26)
$=\sum_{i=1}^{L}\frac{1}{2\sigma^{2}}$ (27)
$=\sum_{i=1}^{L}\frac{\epsilon^{2}}{{8L\log(1/\delta)}}$ (28)
$=\frac{\epsilon^{2}}{{8\log(1/\delta)}}.$ (29)
By Proposition 1.3 of Bun & Steinke (2016), if M provides $\rho$-zCDP, then M is ($\rho+2\sqrt{\rho\log(1/\delta)},\delta)$-DP, $\forall\delta>0$. Let $\epsilon^{\prime}\triangleq\rho+2\sqrt{\rho\log(1/\delta)}$.
Then:
$\epsilon^{\prime}$ $=\rho+2\sqrt{\rho\log(1/\delta)}$ (31)
$\leq\frac{\epsilon^{2}}{{8\log(1/\delta)}}+2\sqrt{\frac{\epsilon^{2}}{{8\log(1/\delta)}}.\log(1/\delta)}$ (32)
$\leq\frac{\epsilon^{2}}{{8\log(1/\delta)}}+\frac{\epsilon}{2}.$ (33)
To satisfy ($\epsilon$, $\delta$)-DP, we need:
$\epsilon\geq\epsilon^{\prime}$ $\impliedby\epsilon\geq\frac{\epsilon^{2}}{{8\log(1/\delta)}}+\frac{\epsilon}{2}$ (35)
$\iff\frac{\epsilon}{2}\geq\frac{\epsilon^{2}}{{8\log(1/\delta)}}$ (36)
$\iff\epsilon\leq 4\log(1/\delta)$ (37)
$\iff\delta\leq\exp{(-\frac{\epsilon}{4})},$ (38)
which is a reasonable assumption, as in practice $\epsilon=O(1)$ and $\delta\ll\frac{1}{d}$ where $d$ is the number of records to protect. In our case, $d=n^{2}$ because we run the privacy-preserving power method on $\bm{A}\in\mathbb{R}^{n\times n}$. ∎
A.3 Proof of Theorem 3.4
Proof.
According to (Hardt & Price, 2014), if $\bm{G}_{l}\sim\mathcal{N}(0,\sigma_{l})^{d\times p}$, then with probability p=99/100 we have:
$\begin{split}&\max_{l=1}^{L}\|\bm{G}_{l}\|\leq\sigma_{l}\cdot\sqrt{d\cdot\log(L)}\,,\\ &\max_{l=1}^{L}\|\bm{U}^{\top}\bm{G}_{l}\|\leq\sigma_{l}\cdot\sqrt{p\cdot\log(L)}\,.\end{split}$ (40)
We can therefore satisfy the noise conditions of Theorem 3.3 with probability $p$ if we choose $\eta=\frac{\sigma_{l}\cdot\sqrt{d\cdot\log(L)}}{\sigma_{k}-\sigma_{q+1}}$, giving us:
$\|(\bm{I}-\bm{X}^{L}{\bm{X}^{L}}^{\top})\cdot\bm{U}_{k}\|\leq\frac{\sigma_{l}\cdot\sqrt{d\cdot\log(L)}}{\sigma_{k}-\sigma_{q+1}}\,.$ (41)
Which leads us to the following Theorem by injecting $\sigma_{l}=\hat{\Delta}_{l}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$.
A.4 Derivation of Approximation 6.2
We provide here the derivation for Approximation 6.2:
Proof.
Let $\bm{M}\in\mathbb{R}^{n\times k}$ with i.i.d $\mathcal{N}(0,1)$-distributed entries. Let $\bm{M}=\bm{Q}\bm{R}$ be its QR factorization (by definition $\bm{i}sQ$ orthogonal and $\bm{R}$ upper triangular). By the Barlett decomposition theorem (Muirhead, 2009), we know that $\bm{Q}$ is a random matrix distributed uniformly in the Stiefel manifold $\mathbb{V}_{k,n}$. Then by Theorem 2.2.1 of Chikuse (2012), we know that a random matrix $\bm{Q}$ uniformly distributed on $\mathbb{V}_{k,n}$ can be expressed as $\bm{Q}=\bm{Z}(\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}}$ with $\bm{Z}$ another matrix with i.i.d $\mathcal{N}(0,1)$-distributed entries. We approximate $\bm{Z}^{\top}\bm{Z}$ as a diagonal matrix and remark that its diagonal elements are distributed as random chi-squared variables with $n$ degrees of liberty. For a matrix $\bm{A}$, we note as $\bm{A}^{\circ n}$ the elementwise Hadamard exponentiation. Then we have:
$\bm{Q}_{ij}^{2}$ $=(\bm{Z}(\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}})_{ij}^{2}$ (42)
$=(\bm{Z}^{\circ 2}((\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}})^{\circ 2})_{ij}$ (43)
$=(\bm{Z}^{\circ 2}(\bm{Z}^{\top}\bm{Z})^{-1})_{ij}.$ (44)
We remark that each element of $\bm{Z}^{\circ 2}$ is distributed as a chi-squared variable with one degree of freedom, therefore $n\cdot\bm{Q}_{ij}^{2}$ is distributed as an i.i.d $F(1,n)$ random variable by definition. For large $n$, we can approximate these $F$ variables as Gaussians with mean $\mu=\frac{n}{n-2}$ and variances $\sigma^{2}=\frac{2n^{2}(n-1)}{(n-2)^{2}(n-4)}$. Therefore $\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{\mu}{n},\frac{\sigma^{2}}{n^{2}})$ and:
$k\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{k\mu}{n},\frac{k^{2}\sigma^{2}}{n^{2}})$ $\sum_{j}\bm{Q}_{ij}^{2}\sim\mathcal{N}(\frac{k\mu}{n},\frac{k\sigma^{2}}{n^{2}}).$ (45)
We can approximate the expectancy of the maximum (noted as m) of $d$ Gaussian variables distributed as $\mathcal{N}(\mu_{2},\sigma_{2}^{2})$ by $m=\sigma_{2}\sqrt{2\cdot\log(d)}+\mu_{2}$ by Lemma 2.3 from Massart (1896). We therefore get approximations of $\mathbb{E}[\max_{ij}k\bm{Q}_{ij}^{2}]$ and $\mathbb{E}[\max_{i}\|\bm{Q}_{i:}\|^{2}]$. We assume that the variances of $\Delta_{l}^{prior}$ and $\hat{\Delta}_{l}$ (as $n$ and $k$ are large) are small and therefore:
$\mathbb{E}[\Delta_{l}^{prior}]^{2}\approx\mathbb{E}[(\Delta_{l}^{prior})^{2}]=\sqrt{\frac{k^{2}\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(kn)}+\frac{k\mu}{n}$ (46)
$\mathbb{E}[\hat{\Delta}_{l}]^{2}\approx\mathbb{E}[(\hat{\Delta}_{l})^{2}]=\sqrt{\frac{k\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(n)}+\frac{k\mu}{n}.$ (47)
Figure 3: Comparison of empirically (in blue) and theoretically (in green) estimated $r(k,n)$ ratios, with $n=12000$ and $k$ ranging between 64 and 4000, with a step of 64. Blue dots represent empirical ratios estimated using $t=5$ runs of the first step of our algorithm. The theoretical ratios are estimated using Theorem 6.2.
A.5 Proof of Theorem 4.1
We provide here the proof for Theorem 4.1: See 4.1
Proof.
First bound: We know by Section 2.4 from Woodruff et al. (2014) that any row $\bm{v}$ of a matrix with orthonormal columns $\bm{Z}$ has squared norm at most 1. Therefore $\max_{i}\|\bm{X}_{i:}\|_{2}\leq 1$ because $\bm{X}$ has orthonormal columns by construction.
Second bound: Without loss of generality and to simplify notation, we note $\bm{X}$ to denote a matrix $\bm{X}_{l}$ computed during the PPM. As proposed in Hardt & Price (2014), we can write any column $x_{c}$ of $\bm{X}$ ($\bm{X}_{:c}$) as $\sum_{i=1}^{n}\alpha_{c}^{i}u_{i}$ where $u_{i}$ denotes the $i$-th eigenvector of $\bm{A}$, and $\alpha_{c}^{i}$ is a scalar. Therefore we have:
$<x_{c},x_{o}>$ $=<\sum_{i=1}^{n}\alpha_{c}^{i}u_{i},\sum_{j=1}^{n}\alpha_{o}^{i}u_{j}>$ (48)
$=\sum_{i=1}^{n}\sum_{j=1}^{n}\alpha_{c}^{i}\alpha_{o}^{i}<u_{i},u_{j}>$ (By bilinearity) (49)
$=\sum_{i=1}^{n}\alpha_{c}^{i}\alpha_{o}^{i}.$ (Orth. columns of $\bm{U}$) (50)
$\bm{X}$ has orthonormal columns, so if $c=o$, then $<x_{c},x_{o}>=\sum_{i=1}^{n}{(\alpha_{c}^{i})}^{2}=1$. Otherwise, we have $<x_{c},x_{o}>=\sum_{i=1}^{n}\alpha_{c}^{i}\alpha_{o}^{i}=0$. We can therefore define a matrix $\bm{B}$ with orthonormal columns such that $\bm{X}=\bm{U}\bm{B}$ and $\bm{B}_{jc}=\alpha_{c}^{j}$. We can therefore write any row $i$ of $\bm{X}$ as $\bm{X}_{i:}=\sum_{j=1}^{p}\bm{U}_{ij}\bm{B}_{j:}$. Therefore
$\|\bm{X}_{i:}\|_{2}$ $=\|\sum_{j=1}^{p}\bm{U}_{ij}\bm{B}_{j:}\|_{2}$ (51)
$\leq\sum_{j=1}^{p}\|\bm{U}_{ij}\bm{B}_{j:}\|_{2}$ (52)
$\leq\sum_{j=1}^{p}|\bm{U}_{ij}|\|\bm{B}_{j:}\|_{2}$ (53)
$\leq\sum_{j=1}^{p}|\bm{U}_{ij}|.$ (Orth. columns of $\bm{B}$) (54)
We then have $\max_{i}\|\bm{X}_{i:}\|_{2}\leq\max_{i}(\sum_{j=1}^{p}|\bm{U}_{ij}|)=\|\bm{U}\|_{\infty}=\mu_{1}(\bm{A})$.
We proposed two bounds for $\|\bm{X}_{i:}\|_{2}$, which are useful in two different cases. We can however note that:
$\|\bm{X}_{i:}\|_{2}\leq 1\text{ and }\|\bm{X}_{i:}\|_{2}\leq\mu_{1}(\bm{A})$ (55)
$\Rightarrow\|\bm{X}_{i:}\|_{2}\leq\min(1,\mu_{1}(\bm{A})),$ (56)
which leads us to Theorem 4.1 giving us a runtime-independent bound.
A.6 Proof of Theorem 5.2
Proof.
It is straightforward to see that if steps 2 and 3 of Algorithm 2 are equivalent to step 1 of Algorithm 1, then the two algorithms are equivalent. Recall that $Y^{(i)}_{\ell}=\bm{A}^{(i)}\bm{X}^{\ell-1}+\bm{G}^{(i)}_{\ell}$ and $\bm{G}^{(i)}_{\ell}\sim\mathcal{N}(0,\Delta_{l}^{2}\nu^{2})^{n\times p}$. Then steps 2 and 3 of Algorithm 2 correspond to:
$\begin{split}\bm{Y}_{\ell}&=SecAgg({\bm{Y}_{\ell}^{(i)}},\{i|1\leq i\leq s\})\\ &=\sum_{i=1}^{s}\bm{Y}_{\ell}^{(i)}\\ &=\sum_{i=1}^{s}(\bm{A}^{(i)}\bm{X}_{\ell-1}+\bm{G}^{(i)}_{\ell})\\ &=\bm{A}\bm{X}^{\ell-1}+\sum_{i=1}^{s}\bm{G}^{(i)}_{\ell}\\ &=\bm{A}\bm{X}^{\ell-1}+\bm{G}_{\ell},\end{split}$ (57)
where $\bm{G}_{\ell}\sim\mathcal{N}(0,\Delta_{l}^{2}\cdot(s\nu^{2}))^{n\times p}$, and we have $s\nu^{2}=\sigma^{2}$ by definition, completing the equivalence proof. ∎
A.7 Application to recommender systems:
Let $s$ be the number of users in our system and $n$ the number of items. Let $\bm{R}\in\mathbb{R}^{s\times n}$ be the user-item interaction matrix, such that $\bm{R}_{ui}=1$ only if user $u$ has interacted with item $i$, and $\bm{R}_{ui}=0$ else. Let $\bm{U}=\text{Diag}(\bm{R}\cdot\bm{1}_{|\mathcal{I}|})$ be the user degrees matrix, and $\bm{I}=\text{Diag}(\bm{1}_{|\mathcal{I}|}^{\top}\cdot\bm{R})$ the item degrees matrix. In Shen et al. (2021); Choi et al. (2023), the normalized interaction matrix is defined as:
$\bm{R}^{\prime}=\bm{U}^{-\frac{1}{2}}\bm{R}\bm{I}^{-\frac{1}{2}},$
and the item-item normalized adjacency matrix as:
$\begin{split}\bm{P}^{\prime}&=\tilde{\bm{R}}^{\top}\tilde{\bm{R}}\\ &=(\bm{U}^{-\frac{1}{2}}\bm{R})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R})\\ &=(\bm{U}^{-\frac{1}{2}}\bm{R}\bm{I}^{-\frac{1}{2}})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R}\bm{I}^{-\frac{1}{2}}).\\ \end{split}$ (58)
To simplify our analysis, we consider $\bm{I}$ public and do not use item-wise normalization in the computation of ideal low pass filter, leaving it for future work. We therefore define $\tilde{\bm{R}}=\bm{U}^{-\frac{1}{2}}\bm{R}$ and $\tilde{\bm{P}}=(\bm{U}^{-\frac{1}{2}}\bm{R})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R})$.
Lemma A.1.
We can use Algorithm 2 with $\Delta_{l}=\sqrt{2}\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ to compute the top-$p$ eigenvectors of $\tilde{\bm{P}}$ in a decentralized setting with under a $(\epsilon,\delta)$-Differential Privacy guarantee.
Proof.
We note $\tilde{\bm{P}}_{ij}$ as the element of $\tilde{\bm{P}}$ at line $i$ and column $j$ and $d_{user}(u)=\sum_{i=0}^{n-1}r_{ui}$. We can write $\tilde{\bm{P}}_{ij}$ as such:
$\begin{split}\tilde{\bm{P}}_{ij}&=((\bm{U}^{-\frac{1}{2}}\bm{R})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R}))_{ij}\\ &=(\bm{U}^{-\frac{1}{2}}\bm{R})^{T})_{i,*}(\bm{U}^{-\frac{1}{2}}\bm{R})_{*,j}\\ &=((\bm{U}^{-\frac{1}{2}}\bm{R})_{*,i})^{\top}(\bm{U}^{-\frac{1}{2}}\bm{R})_{*,j}\\ &=\sum_{u=0}^{s-1}\frac{1}{\sqrt{d_{user}(u)}}r_{ui}\frac{1}{\sqrt{d_{user}(u)}}r_{uj}\\ &=\sum_{u=0}^{s-1}\frac{1}{d_{user}(u)}\cdot r_{ui}\cdot r_{uj}.\\ \end{split}$ (59)
By noticing that $(\bm{R}^{\top}\bm{R})_{ij}=\sum_{u}r_{ui}\cdot r_{uj}$, then we can deduce that $\tilde{\bm{P}}=\sum_{u}\frac{1}{d_{user}(u)}\bm{R}_{u}^{\top}\bm{R}_{u}$. Therefore $\tilde{\bm{P}}$ is partitioned among $s$ users as described in Section 5. Sensitivity: We protect the user at the item-level and use the deletion model of differential privacy to compute the sensitivity of the PPM used with the item-item normalized adjacency matrix ($\bm{A}=\tilde{\bm{P}}$). Therefore we have:
$\bm{A}_{ij}=\sum_{u=0}^{s-1}\frac{1}{d_{user}(u)}\cdot r_{ui}\cdot r_{uj},$ (60)
and
$\bm{A}^{\prime}_{ij}=\sum_{u=0}^{s-1}\frac{1}{d_{user}(u)-1}\cdot r_{ui}^{\prime}\cdot r_{uj}^{\prime},$ (61)
where $r_{ui}^{\prime}=r_{ui}$ except for one user-item interaction, i.e. $r_{vk}=1$ but $r_{vk}^{\prime}=0$. Let $\bm{C}=\bm{A}-\bm{A}^{\prime}$. Let $\mathcal{N}(v)$ be the set of items which user $v$ liked before deletion. We have:
$\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}$ $=\sum_{i}(\sum_{j}|\sum_{u=0}^{s-1}\frac{1}{d_{user}(u)}\cdot r_{ui}\cdot r_{uj}-\sum_{u=0}^{s-1}\frac{1}{d_{user}(u)-1}\cdot r_{ui}^{\prime}\cdot r_{uj}^{\prime}|)^{2}$ (62)
$=\sum_{i}(\sum_{j}|\frac{1}{d_{user}(v)}\cdot r_{vi}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vi}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (63)
$=\sum_{i\in\mathcal{N}(v)}(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vi}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vi}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (64)
$=\sum_{i\in\{\mathcal{N}(v)\backslash k\}}(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vi}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vi}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (65)
$+(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vk}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vk}^{\prime}\cdot r_{vj}^{\prime}|)^{2}.$ (66)
We have:
$\sum_{i\in\{\mathcal{N}(v)\backslash k\}}(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vi}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vi}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (68)
$=\sum_{i\in\{\mathcal{N}(v)\backslash k\}}(\sum_{j\in\{\mathcal{N}(v)\backslash k\}}|\frac{1}{d_{user}(v)}\cdot r_{vi}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vi}^{\prime}\cdot r_{vj}^{\prime}|+\frac{1}{d_{user}(v)}\cdot r_{vi}\cdot r_{vk})^{2}$ (69)
$=(d_{user}(v)-1)\{(d_{user}(v)-1)|\frac{1}{d_{user}(v)}-\frac{1}{d_{user}(v)-1}|+\frac{1}{d_{user}(v)}\}^{2}$ (70)
$=(d_{user}(v)-1)\{|\frac{1}{d_{user}(v)}|+\frac{1}{d_{user}(v)}\}^{2}$ (71)
$=\frac{2(d_{user}(v)-1)}{d_{user}(v)^{2}},$ (72)
and
$(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vk}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vk}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (74)
$=(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vj}|)^{2}$ (75)
$=(d_{user}(v)|\frac{1}{d_{user}(v)}|)^{2}$ (76)
$=1.$ (77)
By noticing that $\frac{2(d_{user}(v)-1)}{d_{user}(v)^{2}}\leq 1$, we can deduce that:
$\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}$ $=\sum_{i\in\{\mathcal{N}(v)\backslash k\}}(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vi}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vi}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (79)
$+(\sum_{j\in\mathcal{N}(v)}|\frac{1}{d_{user}(v)}\cdot r_{vk}\cdot r_{vj}-\frac{1}{d_{user}(v)-1}\cdot r_{vk}^{\prime}\cdot r_{vj}^{\prime}|)^{2}$ (80)
$=\frac{2(d_{user}(v)-1)}{d_{user}(v)^{2}}+1$ (81)
$\leq 2$ (82)
$\implies$ $\sqrt{\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}}\leq\sqrt{2}.$ (83)
By Equation 24,
$\Delta_{l}\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}}$ (84)
$\implies\Delta_{l}\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{2}.$ (85)
A.7.1 Approximation errors comparisons:
As explained in Section 6.1.2, GF-CF and BSPM (with its parameter $T_{b}=1$) use $\bm{U}_{p}$ to compute the ideal low pass filter and filter the interaction matrix $\bm{R}$, yielding $\bm{R}_{p}$. Indeed, we have:
$\bm{R}_{p}=\bm{R}\cdot\bm{I}^{-\frac{1}{2}}\bm{U}_{p}\bm{U}_{p}^{\top}\bm{I}^{\frac{1}{2}}.$ (86)
To study the impact of Differential Privacy on our system, we compute approximations of $\bm{U}_{p}$ (denoted by $\tilde{\bm{U}}_{p}$) using our proposed decentralized PPM, or the PPM versions of Balcan et al. (2016).We have:
$\tilde{\bm{R}}_{p}=\bm{R}\cdot\bm{I}^{-\frac{1}{2}}\tilde{\bm{U}}_{p}\tilde{\bm{U}}_{p}^{\top}\bm{I}^{\frac{1}{2}}.$ (87)
We then define the relative approximation error caused by the use of differential privacy as $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$.
We use $L=3$ to run the Power Method as it is the default hyper-parameter choice from Shen et al. (2021); Choi et al. (2023). We use $p=32$ to have acceptable approximation error (on the order of 0.02-0.05 for numbers bounded between 0 and 1) for reasonable values of $\epsilon$ (5-10). We showed the impact of the Differential Privacy parameter $\epsilon$ on the approximation errors for the EachMovie dataset in Figure 1. We get similar results for the MovieLens and Jester datasets in Figures 5 and  4.
Figure 4: Impact of the Differential Privacy parameter $\epsilon$ on the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to multiple PPM versions, computed with $p=32$ and $L=3$ on the Jester dataset.
Figure 5: Impact of the Differential Privacy parameter $\epsilon$ on the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to multiple PPM versions, computed with $p=32$ and $L=3$ on the MovieLens dataset.